### v1

# Convexity on manifolds without focal points and spectrum of the Laplacian

Abstract.
Convexity on complete Riemannian manifolds has typically been studied under constraints on the sign of sectional curvatures. This is especially so in the study of the effect of convexity on the spectrum of the Laplacian. In this article, we use the language of Jacobi tensors to construct a strictly convex function with bounded gradient on complete simply connected manifolds without focal points having negative Ricci curvature. From this we obtain a lower bound for the growth of the volume of geodesic balls, if the manifold is also assumed to be Kähler. Our primary applications concern the spectrum of the Laplacian. We use the strictly convex function which we construct to show that the spectrum of the Laplacian is absolutely continuous on a class of asymptotically harmonic manifolds. In fact, we completely determine the spectrum of the Laplacian on simply connected noncompact nonflat harmonic manifolds without focal points. For asymptotically harmonic manifolds we also show that equality holds in Cheeger’s inequality for the bottom of the spectrum, where Cheeger’s constant is equal to the constant mean curvature of the horospheres. We show that the well-known description of the spectrum of the Laplacian on Riemannian symmetric spaces of noncompact type and arbitrary rank can also be recovered using these geometric methods.
1. Introduction
Convex functions have long played a fundamental role in Riemannian geometry. By the convexity of a continuous real-valued function on a Riemannian manifold we mean that it is convex in the usual sense along any geodesic. The mere existence or nonexistence of convex functions have geometric and topological implications. Nonconstant convex functions on complete Riemannian manifolds exist only on those with infinite volume [2, Proposition 2.2]. In particular, they don’t exist on compact Riemannian manifolds. There are profound structural results arising from the interaction of curvature and convexity. Under appropriate curvature conditions there are natural geometrically defined convex functions on a Riemannian manifold. On complete simply connected Riemannian manifolds of nonpositive sectional curvature, called Hadamard manifolds, the square of the distance function from a fixed but arbitrary point and the Busemann functions are examples of convex functions which play a crucial role in the geometry of these manifolds. While the former is a smooth convex function the latter is merely a $C^{2}$-convex function [12, §1.6.6]. In the nonnegative sectional curvature situation for instance, a key idea in the fundamental work [8] is the construction of suitable continuous convex functions related to Busemann functions which have appropriate additional properties.
A function which is at least $C^{2}$ is (strictly) convex if its hessian is positive (definite) semi-definite everywhere. If a Riemannian manifold admits a strictly convex function, then it has no compact immersed minimal submanifolds [20, §9.2 C]. In particular, it has no closed geodesics. Such a manifold also allows for a generalization of Cartan’s theorem [29] when this strictly convex function has a minimum: every compact subgroup of the isometry group has a common fixed point. The existence of strict convex functions also has consequences for the $\mathrm{L^{2}}$-spectrum of the Laplacian. The study of the spectrum of the Laplacian on complete Riemannian manifolds has a long, rich history and there are very many results, often with various additional constraints placed on the structure of the manifold and on sectional curvature. For a succint recent overview of some of these aspects, we refer to [9]. We content ourselves by mentioning a few results which use convexity in an essential manner to deduce information on the spectrum to put our work in context, with no pretensions to completeness.
The absolutely continuous part of the spectrum of the Laplacian on complete Riemannian manifolds supporting a sufficiently regular strictly convex function is nontrivial [28, Theorem 2] under certain boundedness conditions on the gradient and the Laplacian of the said function. On complete simply connected manifolds with nonpositive sectional curvature, [28, Theorem 1] exhibited a strictly convex function related to the square of the distance function satisfying the estimates required to ensure the nontriviality of the absolutely continuous spectrum. Moreover, it was also shown that in such a setting there are no eigenvalues in the spectrum. In the case of nonnegative sectional curvature, explicit convex functions have been used to show that there are no eigenfunctions under the additional assumptions on the codimension of the soul of the manifold being $1$ or having a quadratic decay of radial curvatures when the soul is just a point [14, Theorems 4.2, 4.3].
One of the key results of this article is the construction of a $C^{2}$-strictly convex function with bounded gradient on complete simply connected Riemannian manifolds without focal points having negative Ricci curvature (Theorem 3.1). A distinguishing feature here is that we do not make any assumptions on the sign of the sectional curvature. It is well known [18] that there are both compact and noncompact manifolds without focal points whose sectional curvatures change sign. Since the Ricci curvature is an average of sectional curvatures, the sign of the Ricci curvature is a weaker notion when compared to the sign of the sectional curvature. That this is true especially for negative Ricci curvature can be seen from the remarkable result [23, Theorem A] that any smooth manifold of dimension at least three admits a complete Riemannian metric with negative Ricci curvature. Interesting classes of manifolds satisfying our assumptions include Riemannian symmetric spaces of noncompact type as well as simply connected noncompact nonflat harmonic manifolds without focal points. Harmonic manifolds are Einstein [21] i.e, $Ric_{g}=\kappa g$, and have horospheres, which are level sets of Busemann functions, of constant mean curvature $h\geq 0$. If the Ricci curvature $\kappa>0$, then the harmonic manifold is necessarily a compact rank one symmetric space and if $\kappa=0$, then it is necessarily Euclidean. All other simply connected harmonic manifolds have negative Ricci curvature. We remark here that $h>0$ for such manifolds, since a harmonic manifold is Euclidean when $h=0$ [21, §5, Theorem 19].
For the construction of the strictly convex function we make use of Jacobi tensors introduced in [13] and its relation to the Hessian of Busemann functions for manifolds of continuous asymptote. In particular, for manifolds without focal points these Hessians can be seen to be nonnegative. So each Busemann function is convex but not strictly convex since its Hessian vanishes in the direction of the gradient. Motivated by the work done in the context of symmetric spaces of noncompact type in [10], we then average the Busemann functions over the boundary of the geodesic compactification for manifolds without focal points. Here we use [16] for the geometry at infinity of such manifolds. Under the assumption of negative Ricci curvature, a strictly convex $C^{2}$-function with bounded gradient is then obtained. As an immediate consequence, we have that the manifolds under consideration admit smooth strictly convex functions which also has bounded gradient. This follows from the approximation result of Green and Wu [17, Theorem 2(a)]. It is known [26] that the smooth approximation result fails for functions which are merely convex.
Our next application is in the context of manifolds which are, in addition, Kähler. A complex manifold with a complex structure $J$ and a compatible Riemannian metric $g$ is called Kähler if the associated alternating two-form defined by $\omega(X,Y):=g(JX,Y)$ is symplectic. Given any two of these compatible structures the third one can be recovered. ${\mathbb{C}}^{n}$ with the standard metric, bounded domains in ${\mathbb{C}}^{n}$ with the Bergmann metric, Hermitian symmetric spaces with the invariant metric, complex projective spaces with the Fubini-Study metric are some illustrative examples of Kähler manifolds. $C^{2}$-strictly convex functions on Kähler manifolds can be seen to be strictly plurisubharmonic functions. In our context, therefore, we get a smooth strictly plurisubharmonic function with bounded gradient. From this we derive a lower bound for the growth of the volume of metric balls on such manifolds following [24, Lemma 4.2 (i)]. The growth of the volume of the metric balls has been usually studied with lower bounds on Ricci curvature and upper bounds on sectional curvatures [7, §III.4].
As principal applications of Theorem 3.1, we describe the spectrum of the Laplacian for certain classes of manifolds having horospheres of constant mean curvature $h$. Complete simply connected Riemannian manifolds without conjugate points such that horospheres are of constant mean curvature $h\geq 0$ are known as asymptotically harmonic manifolds [21]. On manifolds as in Theorem 3.1, which are in addition asymptotically harmonic, we observe that the strictly convex function that we construct also has bounded Laplacian and proceed to show in Theorem 4.3 that the spectrum is purely absolutely continuous and is contained in $[h^{2}/4,\infty)$. The condition on the bilaplacian required as in [28, Theorem 2] follows automatically since the horospheres have constant mean curvature. In fact, along the way we prove that equality holds in Cheeger’s inequality [6, §4.3, Theorem 3] for asymptotically harmonic manifolds and that Cheeger’s constant is equal to $h$ in Proposition 4.1. Indeed, the proof shows that it is sufficient for horospheres to have constant mean curvature in one direction. Further, we prove in Theorem 4.7 that the spectrum is purely absolutely continuous for simply connected noncompact nonflat harmonic manifolds without focal points and is equal to $[h^{2}/4,\infty)$. These are harmonic manifolds of purely exponential volume growth [21, Theorems 32, 45]. Rank one symmetric spaces of noncompact type and other Damek-Ricci spaces are some examples of this class of manifolds. Since harmonic manifolds are asymptotically harmonic, that the spectrum is purely absolutely continuous and that it is contained in $[h^{2}/4,\infty)$ follow from Theorem 4.3. To prove that the spectrum is equal to $[h^{2}/4,\infty)$, we make use of the Fourier transform developed in this context in [3]. More precisely, using the unitarity of the spherical Fourier transform [3, Theorem 4.7], we can show in the usual way that the multiplication operator $M_{(\lambda^{2}+\frac{h^{2}}{4})}$ is unitarily equivalent to the action of the Laplacian on radial functions to conclude that $[h^{2}/4,\infty)$ is contained in the spectrum.
Higher rank symmetric spaces of noncompact type are not asymptotically harmonic since the horospheres along each direction have constant mean curvature which depend on that direction. Nonetheless, we are able to obtain (Theorem 4.8) that the spectrum of the Laplacian is purely absolutely continuous and is contained in $[\left\lVert\rho^{2}\right\rVert,\infty)$ by making use of the explicit expression for the mean curvature of horospheres and the strictly convex function that we construct in Theorem 3.1. To show that the spectrum is equal to $[\left\lVert\rho^{2}\right\rVert,\infty)$, we use the unitarity of just the spherical Fourier transform as in the case of harmonic manifolds. Note that in contrast to the case of harmonic manifolds where the unitarity of the Fourier transform is not known, it is classical that the Fourier transform in the case of symmetric spaces is unitary. As a consequence, the spectrum can also be completely determined for symmetric spaces via the unitary equivalence between the Laplacian and the multiplication operator $M_{(\left\lVert\lambda\right\rVert^{2}+\left\lVert\rho\right\rVert^{2})}$ given by the Fourier transform. It is worth remarking that in showing that the spectrum of the Laplacian on symmetric spaces of noncompact type is purely absolutely continuous, essential use is made only of the fact that they are manifolds without focal points and have negative Ricci curvature.
2. Preliminaries
Let $M$ be a complete Riemannian manifold and let $\gamma$ be a geodesic in $M$. A vector field $Y$ on $\gamma$ is called a Jacobi field if it satisfies the Jacobi equation
$Y^{{}^{\prime\prime}}+R(\gamma^{{}^{\prime}},Y)\gamma^{{}^{\prime}}=0.$
Here $\gamma^{{}^{\prime}}$ is the tangent vector field of $\gamma$, $R$ is the Riemann curvature tensor on $M$ and $Y^{{}^{\prime}}$ denotes the covariant derivative of $Y$ with respect to $\gamma^{{}^{\prime}}$ along $\gamma$. Two points $p$ and $q$ on $\gamma$ are said to be conjugate along $\gamma$ if there exists a nonzero Jacobi field which vanishes both at $p$ and at $q$. $M$ is called a manifold without conjugate points if for every geodesic $\gamma$ in $M$, no two points on $\gamma$ are conjugate along $\gamma$ or equivalently, if the exponential map is non-singular at every point in $M$. In fact, $M$ has no conjugate points if and only if $\exp_{p}:T_{p}M\to M$ is a covering map for each $p$ in $M$. So if $M$ is also simply connected, then $\exp_{p}$ is a diffeomorphism for all $p\in M$ and hence $M$ is necessarily noncompact. In this case, any two distinct points of $M$ are joined by a unique geodesic. A general reference for the basic Riemannian geometry used in this article is [7].
We now give a brief overview of Jacobi tensors and refer to [13] for a full account. Let $M$ be a complete simply connected Riemannian manifold without conjugate points and let ($SM$) $TM$ denote its (unit) tangent bundle. For $v\in SM$, let $\gamma_{v}$ denote the unit speed geodesic determined by $v$. A normal $(1,1)$-tensor $Y_{v}$ along the geodesic $\gamma_{v}$ is a bundle endomorphism of the normal bundle $N\gamma_{v}$ over $\gamma_{v}$ given by
$Y_{v}:{\mathbb{R}}\to\mathrm{End}N(\gamma_{v})=\bigcup\limits_{t\in{\mathbb{R}}}\mathrm{End}(N(\gamma_{v}(t))),\,N(\gamma_{v}(t)):=\left\{w\in T_{\gamma_{{}_{v}}(t)}M:{\langle w},{\gamma_{v}^{{}^{\prime}}(t)\rangle}_{\gamma_{{}_{v}}(t)}=0\right\}.$
A Jacobi tensor $Y_{v}$ is a normal $(1,1)$-tensor field along a geodesic $\gamma_{v}$ with transversal derivative (i.e. there exists some $t$ such that $ker~{}Y_{v}(t)\cap ker~{}Y_{v}^{{}^{\prime}}(t)=0$) satisfying the Jacobi equation
(2.1) $Y_{v}^{{}^{\prime\prime}}(t)+R(t)Y_{v}(t)=0$
Here $R(t)$ is the symmetric $(1,1)$-tensor along $\gamma_{v}$ induced by the curvature tensor $R$ and defined by
$R(t)w:=R(w,\gamma_{v}^{{}^{\prime}}(t))\gamma_{v}^{{}^{\prime}}(t)$
and $Y_{v}^{{}^{\prime}}$ is the covariant derivative of the tensor $Y_{v}$ along $\gamma_{v}$ with respect to the tangent vector field $\gamma_{v}^{{}^{\prime}}$. For each $s>0$ and $v\in SM$, let $D_{s,v}$ be a solution to (2.1) with the boundary conditions $D_{s,v}(0)=\mathrm{Id}$ and $D_{s,v}(s)=0$. Then $D_{v}:=\lim\limits_{s\to\infty}D_{s,v}$ exists and is referred to as the stable Jacobi tensor $D_{v}$ satisfying $D_{v}(0)=\mathrm{Id}$. A complete simply connected Riemannian manifold without conjugate points is said to be manifold of continuous asymptote if $v\to D_{v}^{{}^{\prime}}(0)$ is a continuous section of the vector bundle $\{(v,\phi):v\in SM,\phi\in\mathrm{End}(N(\gamma_{v})(0))\}$ over $SM$.
Given $t\in[0,\infty)$ and $v\in SM$, let $b_{t,v}:M\to\mathbb{R}$ denote the function defined by
$b_{t,v}(x):=d(x,\gamma_{v}(t))-t$
which, by triangle inequality, is increasing as a function of $t$. Since $b_{t,v}(\cdot)$ is absolutely bounded by $d(\gamma_{v}(0),\cdot)$, the function
$b_{v}:=\lim\limits_{t\to\infty}b_{t,v}$
is defined everywhere on $M$. The function $b_{v}$ is called the Busemann function associated to $v\in SM$. Recall that the level sets of Busemann functions are called horospheres. Horospheres are closed and hence are complete hypersurfaces with the induced metric. Recall that for Hadamard manifolds the Busemann functions are $C^{2}$ [12, §1.6.6]. Eschenburg [13] was able to relax the curvature restriction by making use of convergence properties of Jacobi tensors, and generalized the $C^{2}$-regularity of horospheres to manifolds of continuous asymptote. For such manifolds, [22, §11] gives a compactification known as the Busemann compactification. Let $o\in M$ be an arbitrary but fixed point. Then the Busemann boundary $\partial^{o}M$ is homeomorphic to $S_{p}M$ for any $p\in M$ [22, Proposition 11.4 (2)].
We see that for each $v\in SM$, the Hessian of the Busemann function $b_{v}$ and the stable Jacobi tensor $D_{v}$ are related as follows. The orthogonal projection $x^{\perp}(p,v)$ of $x\in T_{p}M$ onto $(\nabla b_{v}(p))^{\perp}$ will be denoted by $x^{\perp}$ for notational convenience.
Lemma 2.1.
For $p\in M$ and $x\in S_{p}M$, we have ${\mathrm{H}}^{b_{v}}(p)(x,x)={\mathrm{H}}^{b_{v}}(p)(x^{\perp},x^{\perp})$.
Proof.
As $\nabla b_{v}$ is a $C^{1}$ vector field with $\left\lVert\nabla b_{v}\right\rVert=1$, we have that $\nabla_{\nabla b_{v}}\nabla b_{v}=0$. Indeed ${\langle\nabla_{X}\nabla b_{v}},{Y\rangle}={\mathrm{H}}^{b_{v}}(X,Y)$ is symmetric in $X,Y$ and hence for every vector field $X$
${\langle\nabla_{\nabla b_{v}}\nabla b_{v}},{X\rangle}$ $={\langle\nabla_{X}\nabla b_{v}},{\nabla b_{v}\rangle}$
$=\frac{1}{2}X{\langle\nabla b_{v}},{\nabla b_{v}\rangle}=0$
In particular, we see that every integral curve of $\nabla b_{v}$ is a geodesic and hence complete. Then, the orthogonal decomposition $T_{p}M={\mathbb{R}}\nabla b_{v}\oplus(\nabla b_{v})^{\perp}$ enables us to write $x={\langle x},{\nabla b_{v}(p)\rangle}_{T_{p}M}\nabla b_{v}(p)+x^{\perp}$. Using $\nabla_{\nabla b_{v}}\nabla b_{v}=0$ and the bilinearity of ${\mathrm{H}}^{b_{v}}$ we have
${\mathrm{H}}^{b_{v}}(p)(x,x)={\mathrm{H}}^{b_{v}}(p)(x^{\perp},x^{\perp}).$
Notice in the above computation that when $X=\nabla b_{v}$ we have ${\mathrm{H}}^{b_{v}}(\nabla b_{v},\nabla b_{v})=0$ and hence $b_{v}$ is not strictly convex for any $v\in SM$. In fact, to the best of our knowledge, it is not known if Busemann functions are even convex for an arbitrary manifold of continuous asymptote. Using Lemma 2.1 and [13, Proposition 2, Theorem 1] we have the following relation between the stable Jacobi tensors and the Busemann functions in the case of manifolds of continuous asymptote: for $v\in SM$, $p\in M$, and $x\in T_{p}M$,
(2.2) ${\mathrm{H}}^{b_{v}}(p)(x,x)=-{\langle D_{\nabla b_{v}(p)}^{{}^{\prime}}(0)x^{\perp}},{D_{\nabla b_{v}(p)}(0)x^{\perp}\rangle}$
Therefore for asymptotically harmonic manifolds, by taking traces in (2.2), we have $\Delta b_{v}=h=trace(D_{v}^{{}^{\prime}}(0))$ for every $v\in SM$.
3. A strictly convex function on a manifold without focal points
A complete simply connected Riemannian manifold $M$ is said to have no focal points if for every nontrivial Jacobi field $Y$ satisfying $Y(0)=0$, the map $t\mapsto\left\lVert Y(t)\right\rVert$ is strictly increasing for $t>0$ [25, Proposition 4]. In particular, any non-trivial Jacobi field cannot vanish more than once and hence every manifold without focal points is also a manifold without conjugate points. Note that $t\mapsto\left\lVert D_{v}(t)\right\rVert$ is monotonically decreasing for each $v\in SM$. This in particular shows that $M$ is of continuous asymptote [13, §5].
Henceforth in this section we assume that $M$ is a $n$-dimensional complete simply connected Riemannian manifold without focal points. To begin with let us recall the geodesic compactifications of $M$ and its connections to the Riemannian exponential map and the Busemann functions. We refer to [16] for more details on what follows. Given $v\in SM$ there exists a unique geodesic $\gamma_{v}$ in $M$ satisfying $\gamma_{v}^{{}^{\prime}}(0)=v$. Two vectors $v,w\in SM$ are said to be asymptotic if the distance between $\gamma_{v}(t)$ and $\gamma_{w}(t)$ is bounded for all $t\geq 0$. Let $M(\infty)$ be the set of all classes of asymptotic vectors and let $\overline{M}=M\cup M(\infty)$. $\overline{M}$ has a canonical topology with the following property: for any $p\in M$, the exponential map ${\mathrm{Exp}}_{p}:T_{p}M\to M$ extends uniquely to a homeomorphism $\Phi_{p}:\overline{T_{p}M}\to\overline{M}$ where $\overline{T_{p}M}\backslash T_{p}M$ is homeomorphic with the unit sphere $S_{p}M$ in $T_{p}M$. Then for any $p\in M$, there exists a unique geodesic $\gamma_{{}_{\nabla b_{v}(p)}}$ starting at $p$ asymptotic to $\gamma_{v}$. From our discussions in §2, we see that the Busemann boundary $\partial^{o}M$ and the geodesic compactification $M(\infty)$ are homeomorphic.
Further, for each $v\in SM$, the Busemann function $b_{v}$ is convex [13, Theorem 2] i.e., ${\mathrm{H}}^{b_{v}}$ is nonnegative definite for every $v\in SM$. Hence each horosphere is a convex submanifold in $M$. As remarked in the introduction, motivated by the construction on symmetric spaces of noncompact type of a strictly convex function, indicated in [10, penultimate paragraph p. 657], we now construct a strictly convex function out of the Busemann functions, using the geometry at infinity for manifolds without focal points discussed above.
Theorem 3.1.
Let $M$ be a complete non-compact simply connected Riemannian manifold with no focal points and negative Ricci curvature. Then $M$ admits a $C^{2}$-strictly convex function $F$ whose gradient is bounded.
Proof.
Let $o$ be an arbitrary but fixed point in $M$. Let $\{v_{k}\}$ be a countable dense subset in $S_{o}M$. Consider the Dirac measures $\delta_{v_{k}}$ associated to each $v_{k}$. Then ${\eta=\sum\limits_{k=1}^{\infty}\frac{1}{2^{k}}\delta_{v_{k}}}$ defines a probability measure on $S_{o}M$. We identify $S_{o}M$, $M(\infty)$ and $\partial^{o}M$ via the respective homeomorphisms. Let $F(p)=\displaystyle{\int\limits_{M(\infty)}b_{v}(p)~{}\mathrm{d}\eta(v)}$. By dominated convergence theorem, the gradient of $F$ at $p$ is given by $\nabla F(p)=\displaystyle{\int\limits_{M(\infty)}\nabla b_{v}(p)~{}\mathrm{d}\eta(v)}$. Clearly,
$\left\lVert\nabla F(p)\right\rVert=\left\lVert\int\limits_{M(\infty)}\nabla b_{v}(p)~{}\mathrm{d}\eta(v)\right\rVert\leq\int\limits_{M(\infty)}\left\lVert\nabla b_{v}(p)\right\rVert~{}\mathrm{d}\eta(v)=1,\quad\text{as}~{}\left\lVert\nabla b_{v}\right\rVert=1$
i.e., $F$ has bounded gradient. In particular, $F$ is Lipschitz with Lipschitz constant $1$. For $p\in M$ and $w\in S_{p}M$, using (2.2) and that ${\mathrm{H}}^{b_{v}}$ is nonnegative definite we have
$\int\limits_{M(\infty)}{\mathrm{H}}^{b_{v}}(p)(w,w)~{}\mathrm{d}\eta(v)$ $=\left|\int\limits_{M(\infty)}{\langle D_{\nabla b_{v}(p)}^{{}^{\prime}}(0)w},{w\rangle}~{}\mathrm{d}\eta(v)\right|$
$\leq\int\limits_{M(\infty)}|{\langle D_{\nabla b_{v}(p)}^{{}^{\prime}}(0)w},{w\rangle}|~{}\mathrm{d}\eta(v)$
[…]
$\leq\int\limits_{M(\infty)}\left\lVert D_{\nabla b_{v}(p)}^{{}^{\prime}}(0)\right\rVert~{}\mathrm{d}\eta(v)\quad(D_{\nabla b_{v}(p)}^{{}^{\prime}}(0)~{}\text{is a bounded linear map for}~{}v\in S_{p}M).$
The restriction of the continuous map $w\mapsto\left\lVert D_{w}^{{}^{\prime}}(0)\right\rVert$ on $SM$ to $S_{p}M$ shows that there exists some $L>0$ such that $\left\lVert D_{w}^{{}^{\prime}}(0)\right\rVert\leq L$ for all $w\in S_{p}M$. In particular, $\left\lVert D_{\nabla b_{v}(p)}^{{}^{\prime}}(0)\right\rVert\leq L$ for all $v\in S_{o}M$. So another application of the dominated convergence theorem shows that
${\mathrm{H}}^{F}(p)(w,w)=\int\limits_{M(\infty)}{\mathrm{H}}^{b_{v}}(p)(w,w)~{}\mathrm{d}\eta(v).$
The convexity of $F$ follows from the convexity of the Busemann functions. We now claim that $F$ is, in fact, strictly convex. Suppose there exists $p\in M$ and $0\neq w\in T_{p}M$ such that ${\mathrm{H}}^{F}(p)(w,w)=0$ then
$\int\limits_{M(\infty)}{\mathrm{H}}^{b_{v}}(p)(w,w)~{}\mathrm{d}\eta(v)=0$
Since $p\mapsto{\mathrm{H}}^{b_{v}}(p)(w,w)$ is a nonnegative continuous function in $v$, we have ${\mathrm{H}}^{b_{v}}(p)(w,w)=0$ for all $v\in S_{o}M$. Using (2.2),
(3.1) ${\langle D_{\nabla b_{v}(p)}^{{}^{\prime}}(0)w^{\perp}},{D_{\nabla b_{v}(p)}(0)w^{\perp}\rangle}=0,\quad\text{for all}~{}v\in S_{o}M.$
It is known from [13, Section 5] that for $v\in SM$ and all parallel vector fields $x$ normal to $\gamma_{v}$, $\frac{d}{dt}(\left\lVert D_{v}(t)x\right\rVert^{2})\leq 0$ i.e., the function $f_{v,x}(t)=\frac{1}{2}\left\lVert D_{v}(t)x\right\rVert^{2}$ is $C^{2}$ and monotonically decreasing. From (3.1) we see that for each $v\in S_{o}M$, $f_{\nabla b_{v}(p),w^{\perp}}$ has a critical point at 00. We claim that $f_{\nabla b_{v}(p),w^{\perp}}^{{}^{\prime\prime}}(0)=0$. If not, then 00 is either a local minimum or local maximum. Suppose 00 is a local minimum then there exists a $\delta$-neighbourhood of 00 such that $f_{\nabla b_{v}(p),w^{\perp}}(0)\leq f_{\nabla b_{v}(p),w^{\perp}}(t)$ for all $t\in(-\delta,\delta)$. Since $f$ is monotonically decreasing we have $f_{\nabla b_{v}(p),w^{\perp}}(t)=f_{\nabla b_{v}(p),w^{\perp}}(0)$ for all $t\in(0,\delta)$. Hence $f_{\nabla b_{v}(p),w^{\perp}}$ is constant on $(0,\delta)$. Observe that $f_{\nabla b_{v}(p),w^{\perp}}$ is $C^{2}$ and so $f_{\nabla b_{v}(p),w^{\perp}}^{{}^{\prime\prime}}(0)=\lim\limits_{t\to 0_{+}}f_{\nabla b_{v}(p),w^{\perp}}^{{}^{\prime\prime}}(t)=0$, a contradiction. A similar argument could be carried out to show that 00 cannot be a local maximum of $f_{\nabla b_{v}(p),w^{\perp}}$. Hence 00 is a degenerate critical point of $f_{\nabla b_{v}(p),w^{\perp}}$ i.e, $f_{\nabla b_{v}(p),w^{\perp}}^{{}^{\prime\prime}}(0)=0$. Then the Jacobi equation for $D_{\nabla b_{v}(p)}$,
$D_{\nabla b_{v}(p)}^{{}^{\prime\prime}}(0)+R_{\nabla b_{v}(p)}(0)D_{\nabla b_{v}(p)}(0)=0$
gives
00 0 $=f_{\nabla b_{v}(p),w^{\perp}}^{{}^{\prime\prime}}(0)$
$={\langle D_{\nabla b_{v}(p)}^{{}^{\prime\prime}}(0)w^{\perp}},{w^{\perp}\rangle}+\left\lVert D_{\nabla b_{v}(p)}^{{}^{\prime}}(0)w^{\perp}\right\rVert^{2}$
$=-{\langle{R_{\nabla b_{v}(p)}}(0)w^{\perp}},{w^{\perp}\rangle}+\left\lVert D_{\nabla b_{v}(p)}^{{}^{\prime}}(0)w^{\perp}\right\rVert^{2}$
$\implies{\langle R_{\nabla b_{v}(p)}(0)w^{\perp}},{w^{\perp}\rangle}$ $=\left\lVert D_{\nabla b_{v}(p)}^{{}^{\prime}}(0)w^{\perp}\right\rVert^{2}$
(3.2) $\implies K(\nabla b_{v}(p),w^{\perp})$ $=\left\lVert D_{\nabla b_{v}(p)}^{{}^{\prime}}(0)w^{\perp}\right\rVert^{2}\geq 0,\quad\text{for all $v\in S_{o}M$}$
where $K$ denotes the sectional curvature corresponding to the plane spanned by $\nabla b_{v}(p)$ and $w^{\perp}$. Let $x\in S_{p}M$ be such that $x\perp w$. Since $S_{p}M$ is homeomorphic to $M(\infty)$ there exists a unique $v_{o}\in S_{o}M$ such that $\gamma_{v_{o}}$ and $\gamma_{x}$ are asymptotic. As $\{v_{k}\}$ is dense in $S_{o}M$ there exists a subsequence $(v_{k_{m}})$ in $S_{o}M$ converging to $v_{o}$. For each $m$ we have $\nabla b_{v_{{}_{k_{m}}}}(p)\in S_{p}M$ such that $\gamma_{\nabla b_{v_{{}_{k_{m}}}}(p)}$ and $\gamma_{v_{{}_{k_{m}}}}$ are asymptotic and so $v_{{}_{k_{m}}}$ and $\nabla b_{v_{{}_{k_{m}}}}(p)$ correspond to the same element in $M(\infty)$. It then follows that the sequence $(\nabla b_{v_{{}_{k_{m}}}}(p))$ converges to $x$ in $S_{p}M$. We denote the orthogonal projection of $w$ onto $(\nabla b_{v_{{}_{k_{m}}}}(p))^{\perp}$ by $w^{\perp}(m)$. Then
$\left\lVert w-w^{\perp}(m)\right\rVert^{2}$ $=\left\lVert{\langle w},{\nabla b_{v_{k_{m}}}(p)\rangle}\nabla b_{v_{k_{m}}}(p)\right\rVert^{2}$
$=|{\langle w},{\nabla b_{v_{k_{m}}}(p)\rangle}|^{2}\left\lVert\nabla b_{v_{k_{m}}}(p)\right\rVert^{2}$
$=|{\langle w},{\nabla b_{v_{k_{m}}}(p)\rangle}|^{2}$
$\xrightarrow{m\to\infty}0\quad\text{as $\nabla b_{v_{k_{m}}}(p)\to x$ and $x\perp w$}.$
Thus the sequence $(w^{\perp}(m))$ converges to $w$ as $m\to\infty$ and from (3) we have,
$K(\nabla b_{v_{k_{m}}}(p),w^{\perp}(m))\to K(x,w)\geq 0,\quad\text{as $m\to\infty$}$
Now for $\{x_{1},\cdots,x_{n-1}\}$ orthonormal vectors in $T_{p}M$ such that $x_{i}\perp w$, $Ricc(w,w)=\sum\limits_{i=1}^{n}K(x_{i},w)\geq 0$. This is a contradiction to the hypothesis that $M$ has negative Ricci curvature. Hence ${\mathrm{H}}^{F}(p)(w,w)>0$ for every $p\in M$ and every $0\neq w\in T_{p}M$ proving the strict convexity of $F$. ∎
While there are several geometrically motivated convex functions on a Riemannian manifold, typically they have low regularity. Even in situations where such convex functions have higher regularity they often do not have other useful properties such as having bounded gradient. On the one hand, distance functions and Busemann functions are examples of convex functions with bounded gradient having low regularity. On the other hand, on complete simply connected non-compact Riemannian manifolds without focal points, the squared distance function is a smooth strictly convex function on $M$ [11, Lemma 1] whose gradient is not bounded. It is in this context that we have the following approximation result for strictly convex functions with bounded gradient. Note that the smoothing approximation fails if we drop the assumption of strict convexity [26]. Recall that a $C^{2}$-function is strictly convex if its Hessian is positive definite. It was proved in [17, Theorem 2(a)] that a Lipschitz $C^{2}$-function which is strictly convex admits a smooth strictly convex approximation which is also Lipschitz. From this and using the fact that Lipschitz functions have bounded gradient we have the following corollary to our Theorem 3.1.
Corollary 3.2.
Let $M$ be as in Theorem 3.1. Then for any $C>1$, $M$ admits a smooth strictly convex function $f$ such that $\left\lVert\nabla f\right\rVert\leq C$.
The property of having bounded gradient is useful in applications as can be seen below in Corollary 3.3 and Theorem 4.7.
Let $(M,g,J)$ be a Kähler manifold as in the introduction. The complex analogue of convex functions is given by the notion of plurisubharmonic functions. These are defined using the complex Hessian called the Levi form. In local coordinates, the Levi form of a function $f$ is given by
$Lf=4\sum\limits_{i,j=1}^{n}\frac{\partial^{2}f}{\partial z_{i}\partial\overline{z}_{j}}\mathrm{d}z_{i}\otimes\mathrm{d}\overline{z}_{j}$
where $(z_{1},\cdots,z_{n})$ is a local holomorphic coordinate system. Let $(x_{1},y_{1},\cdots,x_{n},y_{n})$ be the corresponding normal coordinates with $z_{k}=x_{k}+iy_{k}$, $k=1,\cdots,n$. A real-valued function on a Kähler manifold is said to be (strictly) plurisubharmonic, if the Levi form is positive (definite) semi-definite. Since $M$ is Kähler, the Levi form $L_{f}$ of $f$ is expressed in terms of Hessian of $f$ using the real coordinates as
$L_{f}={\mathrm{H}}^{f}\left(\frac{\partial}{\partial x_{1}},\frac{\partial}{\partial x_{1}}\right)+{\mathrm{H}}^{f}\left(\frac{\partial}{\partial y_{1}},\frac{\partial}{\partial y_{1}}\right).$
Therefore if $f$ is a strictly convex $C^{2}$-function, then $f$ is strictly plurisubharmonic. As an application to Corollary 3.2, using [24, Lemma 4.2 (i)], we show the following volume growth comparison for metric balls.
Corollary 3.3.
If $M$ as in Theorem 3.1 is also Kähler of (complex) dimension $m$, then $M$ admits a smooth strictly plurisubharmonic function with bounded gradient. Moreover, for each $p\in M$ there exists a constant $C_{2}(p)>0$ such that ,
$V_{p}(r)\geq C_{2}(p)r^{m},\quad r\geq 1$
where $V_{p}(r)$ is the volume of a metric ball of radius $r$ about $p$ in $M$.
4. Spectral theoretic applications
We begin with a sketch of the basic functional analytic setup concerning the spectrum of the Laplacian following [4]. Let $M$ be a complete Riemannian manifold. Then the (positive) Laplacian $\Delta$ is an essentially self-adjoint operator with the space of compactly supported smooth functions on $M$ as the domain. We also denote its unique self-adjoint extension, having the domain $\mathcal{D}(\Delta):=\{f\in L^{2}(M):\Delta f\in L^{2}(M)\}$, by $\Delta$. Using the functional calculus for $\Delta$ we have the family of spectral projection operators $\prod_{E}:=\chi_{E}(\Delta)$ for each Borel subset $E$ of ${\mathbb{R}}$ referred to as the resolution of $\Delta$. The support of these projections and the spectrum of $\Delta$ are related in the following way. We say that $\lambda\in{\mathbb{C}}$ is in the spectrum $\sigma(\Delta)$ of $\Delta$ if and only if $\prod_{(\lambda-\epsilon,\lambda+\epsilon)}\neq 0$ for all $\epsilon>0$. If $\prod_{\{\lambda\}}\neq 0$, then $\lambda$ is an eigenvalue of $\Delta$ and the range of $\prod_{\{\lambda\}}$ is the corresponding eigenspace. The spectral theorem gives a unitary equivalence between $\Delta$ and a multiplication operator on a suitable Hilbert space of the form $\mathrm{L}^{2}(X,\mu)$, where $\mu$ is a $\sigma-$ finite Borel measure on $X$. By the Lebesgue decomposition, $\mu$ decomposes as $\mu=\mu_{pp}+\mu_{ac}+\mu_{sc}$ into pure point measure, absolutely continuous and singularly continuous measures with respect to the Lebesgue measure, and consequently we have that
(4.1) $\mathrm{L}^{2}(X,\mu)=\mathrm{L}^{2}(X,\mu_{pp})\oplus\mathrm{L}^{2}(X,\mu_{ac})\oplus\mathrm{L}^{2}(X,\mu_{sc}).$
This induces a decomposition $\mathrm{L}^{2}(M)=\mathcal{H}_{pp}\oplus\mathcal{H}_{ac}\oplus\mathcal{H}_{sc}$ from which we have the associated decomposition of the spectrum of $\Delta$ namely, $\sigma(\Delta)=\overline{\sigma_{pt}(\Delta)}\cup\sigma_{ac}(\Delta)\cup\sigma_{sc}(\Delta)$ where
$\sigma_{ac}(\Delta)=\sigma(\Delta_{|_{\mathcal{H}_{ac}}}),\quad\sigma_{sc}(\Delta)=\sigma(\Delta_{|_{\mathcal{H}_{sc}}}),\quad\overline{\sigma_{pt}(\Delta)}=\sigma(\Delta_{|_{\mathcal{H}_{pp}}})$
Here we take the closure for the set of all eigenvalues as it is not necessarily closed. We refer to $\sigma_{ac}(\Delta)$ and $\sigma_{sc}(\Delta)$ as the absolutely continuous and singularly continuous parts of $\sigma(\Delta)$ respectively and their union as the continuous part $\sigma_{cont}(\Delta)$ of $\sigma(\Delta)$.
In order to analyse the nature of the spectrum of the Laplacian, we invoke [28, Theorem 2]. This result demands boundedness of the gradient and of the Laplacian of a strictly convex function as well as the additional constraint of nonpositivity on the bilaplacian and in turn yields the absolute continuity of the spectrum. Consequently, it enables us to exclude the existence of the point spectrum and the singularly continuous spectrum. As pointed out in [28], the seemingly unnatural additional condition that $\Delta^{2}F\leq 0$ is, in fact, necessary. Indeed, if this condition on the bilaplacian fails, Hadamard manifolds with eigenvalues in the spectrum are constructed there. The proof of absolute continuity of the spectrum is then derived from an estimate implementing Kato’s theory of $H-$smooth operators in Hilbert spaces. We shall observe below that the condition on the bilaplacian is satisfied for $M$ as in Theorem 3.1 which are in addition asymptotically harmonic.
We first briefly discuss some facts regarding the bottom $\lambda_{0}(M)$ of the spectrum for asymptotically harmonic manifolds. For a complete noncompact Riemannian manifold $M$ it is characterized by [6, §4.4]
(4.2) $\lambda_{0}(M)=\inf\limits_{f\neq 0}\frac{\int\limits_{M}|\nabla f|^{2}}{\int\limits_{M}|f|^{2}}$
where $f$ varies over the completion $\mathfrak{H}(M)$ of $C_{c}^{\infty}(M)$ with respect to the norm $\left\lVert f\right\rVert=\left\lVert f\right\rVert_{\rm{L}^{2}(M)}+\left\lVert\nabla f\right\rVert_{\rm{L}^{2}(M)}$. Let $|\Omega|$ denote the volume of a submanifold $\Omega$ in $M$. Then the Cheeger’s constant of $M$ is given by [6, §4.3]
$h_{Cheeger}(M):=\inf\limits_{\Omega\subset M}\frac{|\partial\Omega|}{|\Omega|}$
where the infimum varies over all open, connected submanifolds $\Omega$ of $M$ with compact closure and smooth boundary $\partial\Omega$. The importance of this global geometric invariant is illustrated by the following fundamental inequality relating it to the global analytic invariant $\lambda_{0}(M)$ [6, §4.3, Theorem 3],
(4.3) $\frac{h_{Cheeger}^{2}(M)}{4}\leq\lambda_{0}(M).$
Proposition 4.1.
Let $M$ be an asymptotically harmonic manifold with horospheres of mean curvature $h$ i.e. $\Delta b_{v}=h$ for all $v\in SM$. Then we have $h_{Cheeger}=h$ and $\lambda_{0}(M)=h^{2}/4$.
Proof.
For an open connected $\Omega$ in $M$ with compact closure and smooth boundary $\partial\Omega$ we have, by Stokes’ theorem, that
$h|\Omega|=\int\limits_{\Omega}\Delta b_{v}=\left|\int\limits_{\partial\Omega}{\langle\nabla b_{v}},{\nu\rangle}\right|\leq|\partial\Omega|,\quad(\text{since$\left\lVert\nabla b_{v}\right\rVert=1$})$
where $\nu$ is the outer unit normal vector field of $\Omega$ along $\partial\Omega$. Thus,
(4.4) $h\leq\inf\limits_{\Omega\subseteq M}\frac{|\partial\Omega|}{|\Omega|}=h_{Cheeger}$
where the infimum varies over $\Omega$ as above in $M$. Let $o\in M$ be arbitrary. Since $M$ is simply connected without conjugate points the exponential at $o$ is a diffeomorphism. Let $d(o,\cdot):M\to{\mathbb{R}}$ denote the distance from $o$. By taking the sequence of functions $f_{n}(x)=e^{-\frac{1}{2}(h+\frac{1}{n})d(o,x)}$ in $\mathfrak{H}(M)$, $n\in{\mathbb{N}}$, we see from (4.2) that
(4.5) $\lambda_{0}(M)\leq\frac{h^{2}}{4}$
Combining (4.3), (4.4), (4.5) we have
(4.6) $h=h_{Cheeger}\quad\text{and}\quad\lambda_{0}(M)=\frac{h^{2}}{4}.$
We remark here that for the proposition above to hold it suffices that there is a $v\in SM$ such that the corresponding horospheres have constant mean curvature.
Lemma 4.2.
Let $M$ and $F$ be as in Theorem 3.1. Further, assume that $M$ is asymptotically harmonic with $h$ as the constant mean curvature of the horospheres. Then $F$ has bounded Laplacian and $\Delta^{2}F=0$.
Proof.
Since $\Delta$ is the trace of Hessian, by choosing an orthonormal basis $\{E_{i}(p)\}_{i=1}^{n}$ of $T_{p}M$ we have
$\sum\limits_{i=1}^{n}{\mathrm{H}}^{F}(p)(E_{i}(p),E_{i}(p))$ $=\sum\limits_{i=1}^{n}\int\limits_{M(\infty)}{\mathrm{H}}^{b_{v}}(p)(E_{i}(p),E_{i}(p))~{}\mathrm{d}\eta(v)$
$=\int\limits_{M(\infty)}\sum\limits_{i=1}^{n}{\mathrm{H}}^{b_{v}}(p)(E_{i}(p),E_{i}(p))~{}\mathrm{d}\eta(v)$
$\Delta F(p)$ $=\int\limits_{M(\infty)}\Delta b_{v}(p)~{}\mathrm{d}\eta(v)=h$
and the lemma follows. ∎
Theorem 4.3.
If $M$ as in Theorem 3.1 is also asymptotically harmonic with $h$ as the constant mean curvature of the horospheres, then the spectrum of the Laplacian is purely absolutely continuous and is contained in $[\frac{h^{2}}{4},\infty)$.
Proof.
The absolute continuity of the spectrum follows by combining [28, Theorem 2] with Lemma 4.2. That the spectrum of Laplacian is contained in $[\frac{h^{2}}{4},\infty)$ follows from (4.6). ∎
Next we briefly describe harmonic manifolds, only mentioning aspects which are needed in the sequel and to set notation, and we refer to [1, Chapter $6$], and to the comprehensive exposition [21] for the details. A function $f$ on a Riemannian manifold $(M,g)$ is said to be radial around a point $o$ if $f$ is constant on geodesic spheres $S_{o}(r)$ of radius $r$ centered at $o$. $M$ is said to be harmonic if the volume density in normal coordinates centered at $o$ is a radial function. Given a continuous function $f$, its radialisation about $o$ given by
$R_{o}(f)(z):=\int\limits_{S_{o}(r)}f(y){\rm{d}}\sigma^{r}(y)$
where $\sigma^{r}$ denotes surface area measure on $S_{o}(r)$ (induced from the Riemannian metric on $M$), normalized to have mass one. Note that for $f$ radial around $o$, we have that $R_{o}f=f$. Harmonicity of the manifold can be seen to be equivalent to the fact that the radialisation operator $R_{o}$ commutes with the Laplacian $\Delta$. Harmonic manifolds are necessarily Einstein manifolds with Ricci curvature $Ric(g)=\kappa g$. Recall that we have the following description of simply connected harmonic manifolds such that $\kappa\geq 0$:
  • (a)
    If $\kappa>0$, then $M$ is a compact rank one symmetric space
  • (b)
    If $\kappa=0$, then $M$ is Euclidean.
Thus simply connected noncompact nonflat harmonic manifolds necessarily have constant negative Ricci curvature. For every $v\in SM$, the Busemann function $b_{v}$ satisfies $\Delta b_{v}=h$ (see for instance [5, §2, p. 33]) and so it follows that every harmonic manifold is also asymptotically harmonic. Consequently we obtain the following corollary to Theorem 4.3.
Corollary 4.4.
Let $M$ be a complete simply connected noncompact nonflat harmonic manifold $M$ without focal points. Then the spectrum of the Laplacian is purely absolutely continuous and is contained in $[h^{2}/4,\infty)$.
We now proceed to show that the spectrum is actually equal to the interval $[h^{2}/4,\infty)$, by making use of the unitarity of the spherical Fourier transform for radial functions on harmonic manifolds of purely exponential volume growth studied in [3]. Notice that in [3] unitarity is shown only in dimensions greater than $5$. However, the only simply connected noncompact nonflat harmonic manifolds of dimension less than equal to $5$ are Riemannian symmetric spaces of noncompact type. This case is handled in Theorem 4.8 below. First, we observe from [21, Theorems 32, 45] that, noncompact harmonic manifolds without focal points fall into the class of noncompact harmonic manifolds of purely exponential volume growth. Let $o$ be an arbitrary but fixed point in $M$. Then for $f\in C_{c}^{\infty}(M)$ which is radial about $o$, the spherical Fourier transform ${\mathcal{F}}^{o}f$ is given by
${\mathcal{F}}^{o}(f)(\lambda,v)=\int\limits_{M}f(y)(\phi_{\lambda}\circ d_{o})(y){\rm{d}}y$
for $\lambda\in{\mathbb{C}}$ and $v\in S_{o}M$. Here $\phi_{\lambda}$ is the unique function on $[0,\infty)$ with $\phi_{\lambda}(0)=1$ and satisfying $\Delta(\phi_{\lambda}\circ d_{o})=(\lambda^{2}+\frac{h^{2}}{4})(\phi_{\lambda}\circ d_{o})$, where $d_{o}$ denotes the distance from $o$. Let ${\rm{L}}_{o}^{2}(M)$ denote the closed subspace of ${\rm{L}}^{2}(M)$ consisting of those functions which are radial around $o$ in ${\rm{L}}^{2}(M)$. It is known from [3, Theorem 4.7] that the ${\mathcal{F}}^{o}$ is an unitary operator from ${\rm{L}}_{o}^{2}(M)$ onto ${\rm{L}}^{2}([0,\infty),C_{0}|c(\lambda)|^{-2}{\rm{d}}\lambda)$. Here the $c$-function coincides with the Harish Chandra $c$-function on $Im(\lambda)\leq 0$, for rank one symmetric spaces of noncompact type.
We recall that for a real-valued measurable function $f$ on a $\sigma-$finite measure space $(X,\mu)$, the multiplication operator $M_{f}$ on ${\rm{L}}^{2}(X,\mu)$ is self-adjoint with domain $\mathcal{D}(M_{f}):=\{g\in{\rm{L}}^{2}(X,\mu):M_{f}g\in{\rm{L}}^{2}(X,\mu)\}$. Let $m$ be the function on $[0,\infty)$ defined by $m(\lambda)=\lambda^{2}+\frac{h^{2}}{4}$. Then we see that the multiplication operator $M_{m}$ on ${\rm{L}}^{2}([0,\infty),C_{0}|c(\lambda)|^{-2}{\rm{d}}\lambda)$ is self-adjoint on
$\mathcal{D}(M_{m})=\{f\in{\rm{L}}^{2}([0,\infty),C_{0}|c(\lambda)|^{-2}{\rm{d}}\lambda):M_{m}f\in{\rm{L}}^{2}([0,\infty),C_{0}|c(\lambda)|^{-2}{\rm{d}}\lambda)\}.$
Moreover, the spectrum of a multiplication operator $M_{m}$ is equal to the essential range of $m$ [4, Theorem 4.5]. Here
$ess-range(m):=\{x\in{\mathbb{R}}:\mu(m^{-1}(B(x,\epsilon)))>0~{}\forall~{}\epsilon>0\}.$
Now to determine the spectrum of $M_{m}$ we argue as follows. For any $x\in{\mathbb{R}}$ and $\epsilon>0$,
$m^{-1}(B(x,\epsilon))=\{\lambda\in[0,\infty):x-\epsilon<\lambda^{2}+\frac{h^{2}}{4}<x+\epsilon\}$
Note that when $x<\frac{h^{2}}{4}$ and $\epsilon=-(x-\frac{h^{2}}{4})$, $m^{-1}(B(x,\epsilon))$ is empty and so $x\notin ess-range(m)$. Equivalently, $ess-range(m)\subseteq[\frac{h^{2}}{4},\infty)$. Now for any $x\in[\frac{h^{2}}{4},\infty)$ and $\epsilon>0$ we have,
$m^{-1}(B(x,\epsilon))=\begin{cases}\left[0,\sqrt{x-\frac{h^{2}}{4}+\epsilon}\right)&\text{if}~{}x-\frac{h^{2}}{4}-\epsilon<0\\ \\ \left[\sqrt{x-\frac{h^{2}}{4}-\epsilon},\sqrt{x-\frac{h^{2}}{4}+\epsilon}\right)&\text{if}~{}x-\frac{h^{2}}{4}-\epsilon\geq 0.\\ \end{cases}$
Using the estimates on $|c(\lambda)|^{-1}$ given in [3, §4.1] we see that, for $C>0$ and $K=\sqrt{x-\frac{h^{2}}{4}+\epsilon}>0$,
$\int\limits_{m^{-1}(B(x,\epsilon))}C_{0}|c(\lambda)|^{-2}{\rm{d}}\lambda\geq\frac{C_{0}}{C}\int\limits_{m^{-1}(B(x,\epsilon))}|\lambda|^{2}{\rm{d}}\lambda>0.$
It follows that $ess-range(m)=[\frac{h^{2}}{4},\infty)$ and so the spectrum of $M_{m}$ is equal to $[\frac{h^{2}}{4},\infty)$. For functions that are not necessarily radial, a notion of Fourier transform, relative to a point, was introduced in [3, §5] where it was shown to be an isometry but the unitarity is not known. This Fourier transform coincides with the spherical Fourier transform ${\mathcal{F}}^{o}$ for radial functions [3, Proposition 5.3]. Then by using [5, Lemma 6.5], we see that
${\mathcal{F}}^{o}\Delta f=M_{m}{\mathcal{F}}^{o}f,\quad f\in{\rm{L}}_{o}^{2}(M).$
Since ${\mathcal{F}}^{o}$ is unitary and $\mathcal{D}(M_{m})={\mathcal{F}}^{o}(\mathcal{D}(\Delta)\cap{\rm{L}}_{o}^{2}(M))$, we have the following lemma.
Lemma 4.5.
The spectrum of the Laplacian $\Delta$ on $\mathcal{D}(\Delta)\cap{\rm{L}}_{o}^{2}(M)$ is equal to $[h^{2}/4,\infty)$.
We now claim that,
Lemma 4.6.
$[h^{2}/4,\infty)\subseteq\sigma(\Delta)$ where $\sigma(\Delta)$ is the spectrum of $\Delta$ with domain $\mathcal{D}(\Delta)$.
Proof.
If $\mu$ is not in $\sigma(\Delta)$, then $\Delta-\mu:\mathcal{D}(\Delta)\to{\rm{L}}^{2}(M)$ is invertible with bounded inverse. In particular, $\Delta-\mu$ is injective on $\mathcal{D}(\Delta)\cap{\rm{L}}_{o}^{2}(M)$. If $f\in{\rm{L}}_{o}^{2}(M)$, then there exists a unique $g$ in $\mathcal{D}(\Delta)$ such that $(\Delta-\mu)g=f$. We have
$(\Delta-\mu)g=f=R_{o}f=R_{o}(\Delta-\mu)g=(\Delta-\mu)R_{o}g$
and by uniqueness $R_{o}g=g$ i.e., $\Delta-\mu$ is surjective from $\mathcal{D}(\Delta)\cap{\rm{L}}_{o}^{2}(M)$ onto ${\rm{L}}_{o}^{2}(M)$ and hence invertible. Since $(\Delta-\mu)^{-1}$ is bounded on ${\rm{L}}^{2}(M)$ and ${\rm{L}}_{o}^{2}(M)$ is a closed subspace, $(\Delta-\mu)^{-1}$ is also bounded on $\mathcal{D}(\Delta)\cap{\rm{L}}_{o}^{2}(M)$. The claim now follows from Lemma 4.5. ∎
The equality of the spectrum now follows from Corollary 4.4 and Lemma 4.6. Therefore, we have the following theorem.
Theorem 4.7.
Let $M$ be a simply connected noncompact nonflat harmonic manifold without focal points. Then the spectrum of Laplacian is purely absolutely continuous and is equal to $[h^{2}/4,\infty)$.
[…]
$R_{v}e_{i}=-R(v,e_{i})v=\lambda_{i}(v)e_{i}.$
From the structure theory for noncompact connected real semisimple Lie algebras, we have the set of all positive roots denoted by $\sum^{+}$. Let $m_{\alpha}$ denote the multiplicity of $\alpha\in\sum^{+}$. The finite reflection group $W$ generated by the roots is known as the Weyl group. We denote the tangent vector that is dual to $2\rho=\sum\limits_{\alpha\in\sum^{+}}m_{\alpha}\alpha$ also by $2\rho$. The arguments in [27, Lemma 1] show that for each $v\in S_{o}M$, $\Delta b_{v}$ is the constant function given by $\sum\limits_{i}\sqrt{\lambda_{i}(v)}={\langle v},{2\rho\rangle}$. This constant is the mean curvature of horospheres with respect to $b_{v}$. Here we argue as in Proposition 4.1 by using Stoke’s theorem to get a lower bound for Cheeger’s constant namely, ${\langle v},{2\rho\rangle}\leq h_{Cheeger}$ for all $v\in S_{o}M$. In particular, for $v=\frac{\rho}{\left\lVert\rho\right\rVert}$ we have $\left\lVert\rho\right\rVert\leq h_{Cheeger}$ and hence by Cheeger’s inequality we obtain
(4.7) $\left\lVert\rho\right\rVert^{2}\leq\lambda_{0}(M).$
Observe that $M$ has negative Ricci curvature [12, §2.14, p. 96] and has no focal points. With notations as in Theorem 3.1 we have,
  1. (i)
    $\left\lVert\nabla F\right\rVert\leq 1$ since $\left\lVert\nabla b_{v}\right\rVert=1$ for all $v\in S_{o}M$.
  2. (ii)
    From $\Delta b_{v}={\langle v},{2\rho\rangle}$ for all $v\in S_{o}M$, we have $|\Delta b_{v}|\leq\left\lVert 2\rho\right\rVert$. Thus $\Delta b_{v}$ is bounded independent of $v\in S_{o}M$ by $\left\lVert 2\rho\right\rVert$ and so $|\Delta F|\leq\left\lVert 2\rho\right\rVert$.
  3. (iii)
    For each $v\in S_{o}M$, $\Delta b_{v}$ is the constant function ${\langle v},{2\rho\rangle}$ and so we have $\Delta^{2}b_{v}=0$. It follows that, $\Delta^{2}F=0$.
The absolute continuity of the spectrum $\sigma(\Delta)$ is then established by invoking [28, Theorem 2]. Further, using (4.7) we have,
Theorem 4.8.
Let $M$ be a symmetric space of noncompact type. Then the spectrum of the Laplacian on $M$ is absolutely continuous and is equal to $[\left\lVert\rho^{2}\right\rVert,\infty)$.
Here we use the spherical Fourier transform for the $K$-bi-invariant functions on $G$ to prove the equality of the spectrum. The radial part of the Laplacian $\Delta$ for the action of $K$ on $G/K$ is given by [19, §4.5, p. 427],
(4.8) $\Delta_{rad}\phi_{\lambda}=({\langle\lambda},{\lambda\rangle}+{\langle\rho},{\rho\rangle})\phi_{\lambda},\quad\lambda\in\mathfrak{a}_{{\mathbb{C}}}^{\ast}$
where $\mathfrak{a}_{{\mathbb{C}}}^{\ast}$ the complexification of the dual $\mathfrak{a}^{\ast}$ of $\mathfrak{a}$ and $\phi_{\lambda}(gK)=\displaystyle{\int\limits_{K}e^{(i\lambda+\rho)A(kg)}~{}{\rm{d}}k}$ is the unique spherical function on $G$ with the normalization $\phi_{\lambda}(eK)=1$ and $A$ is referred to as the Iwasawa projection $A:G\to{\mathfrak{a}}$. Two spherical functions are identical up to the action by the Weyl group $W$ on $\mathfrak{a}_{{\mathbb{C}}}^{\ast}$. Let us denote the space of smooth compactly supported $K-$invariant functions on $G/K$ by $C_{c}^{\infty}(K\backslash G/K)$ and its closure in ${\rm{L}}^{2}(G/K)$ by ${\rm{L}}_{rad}^{2}(G/K)$. Using the inversion formula for the spherical transform [19, Chapter 4, Theorem 7.5] the map $f\mapsto\widehat{f}$
$\widehat{f}(\lambda)=\int\limits_{G/K}f(gK)\phi_{\lambda}(gK),\quad f\in C_{c}^{\infty}(K\backslash G/K)$
[…]
(4.9) $\int\limits_{m^{-1}(B(x,\epsilon))}|\textbf{c}(\lambda)|^{-2}{\rm{d}}\lambda>0.$
for $x\geq\left\lVert\rho\right\rVert^{2}$ and $\epsilon>0$. Note that $|\textbf{c}(\lambda)|^{2}$ defined on ${\mathfrak{a^{\ast}}}$ is invariant under the action of $W$ i.e., $|\textbf{c}(w\lambda)|^{2}=|\textbf{c}(\lambda)|^{2}$ for all $w\in W$ [19, §IV.7, Equation 16] and hence descends to a function on ${\mathfrak{a}^{\ast}}/W$. Since $\lambda\mapsto\left\lVert\lambda\right\rVert^{2}$ is continuous and
$m^{-1}(B(x,\epsilon))=\left\{\lambda\in{\mathfrak{a}^{\ast}}/W:x-\epsilon-\left\lVert\rho\right\rVert^{2}<\left\lVert\lambda\right\rVert^{2}<x+\epsilon-\left\lVert\rho^{2}\right\rVert\right\}$
we see that $m^{-1}(B(x,\epsilon))$ contains an open ball in $\mathfrak{a}^{\ast}/W$. Using [15, Chapter 3, p. 101, Theorem 6.3.4], we have that the map $\lambda\mapsto|\textbf{c}(\lambda)|^{-2}$ is analytic on $\mathfrak{a}^{\ast}$ and thereby analytic on $\mathfrak{a}^{\ast}/W$. Thus, (4.9) holds for $x\geq\left\lVert\rho\right\rVert^{2}$ and $\epsilon>0$. This, in particular, shows that the spectrum of $\Delta_{rad}$ is equal to $[\left\lVert\rho\right\rVert^{2},\infty)$. We argue as in Lemma 4.6 to show the equality in Theorem 4.8.
5. Funding
This work was supported by the Centre for Operator Algebras, Geometry, Matter and Spacetime, Ministry of Education, Government of India through Indian Institute of Technology Madras [SB22231267MAETWO008573 to A.P. and S.B.]; and by the National Board for Higher Mathematics, Department of Atomic Energy, Government of India [Ref.No:0203/11/2019-R&D-II to S.B.]

### v2

@@
-Convexity on manifolds without focal points and spectrum of the Laplacian
+Convexity on manifolds without focal points and applications
 Abstract.
-Convexity on complete Riemannian manifolds has typically been studied under constraints on the sign of sectional curvatures.
-This is especially so in the study of the effect of convexity on the spectrum of the Laplacian.
-In this article, we use the language of Jacobi tensors to construct a strictly convex function with bounded gradient on complete simply connected manifolds without focal points having negative Ricci curvature.
-From this we obtain a lower bound for the growth of the volume of geodesic balls, if the manifold is also assumed to be Kähler.
+In this article, we study strict convexity on simply connected Riemannian manifolds without focal points, a class of manifolds which contains not only Hadamard manifolds but also examples of manifolds whose sectional curvatures change sign.
+Using the geometrically defined convex functions on such manifolds, namely the distance function and the Busemann function, strictly convex functions with various desired additional properties are constructed.
+From the former, we derive interesting consequences such as the continuity of the isoperimetric profile function without conditions on the sectional curvature and, if the manifolds are also Kähler, Steinness as well as a lower bound on the volume growth of metric balls.
 Our primary applications concern the spectrum of the Laplacian.
-We use the strictly convex function which we construct to show that the spectrum of the Laplacian is absolutely continuous on a class of asymptotically harmonic manifolds.
-In fact, we completely determine the spectrum of the Laplacian on simply connected noncompact nonflat harmonic manifolds without focal points.
-For asymptotically harmonic manifolds we also show that equality holds in Cheeger’s inequality for the bottom of the spectrum, where Cheeger’s constant is equal to the constant mean curvature of the horospheres.
-We show that the well-known description of the spectrum of the Laplacian on Riemannian symmetric spaces of noncompact type and arbitrary rank can also be recovered using these geometric methods.
+We prove that the absolutely continuous part of the spectrum contains a certain infinite interval assuming only the existence of a point with respect to which the radial curvatures are nonpositive.
+This is a substantial generalisation of the corresponding result for Hadamard manifolds.
+In the latter case of Busemann functions, we use the geometry at infinity to give a new construction of a strictly convex function assuming that the Ricci curvature is negative.
+We then apply this to show that the spectrum is purely absolutely continuous on a class of manifolds for which all the horospheres have the same constant mean curvature.
+In particular, for a well-studied subclass for which the condition on Ricci curvature is automatically satisfied, we completely determine the spectrum.
+This construction also enables us to recover the classical result on the spectrum in the case of symmetric spaces of noncompact type and higher rank even though the mean curvatures of horospheres there are not constant.
 1. Introduction
@@
 The mere existence or nonexistence of convex functions have geometric and topological implications.
-Nonconstant convex functions on complete Riemannian manifolds exist only on those with infinite volume [2, Proposition 2.2].
+Nonconstant convex functions on complete Riemannian manifolds exist only on those with infinite volume.
 In particular, they don’t exist on compact Riemannian manifolds.
@@
 Under appropriate curvature conditions there are natural geometrically defined convex functions on a Riemannian manifold.
-On complete simply connected Riemannian manifolds of nonpositive sectional curvature, called Hadamard manifolds, the square of the distance function from a fixed but arbitrary point and the Busemann functions are examples of convex functions which play a crucial role in the geometry of these manifolds.
-While the former is a smooth convex function the latter is merely a $C^{2}$-convex function [12, §1.6.6].
-In the nonnegative sectional curvature situation for instance, a key idea in the fundamental work [8] is the construction of suitable continuous convex functions related to Busemann functions which have appropriate additional properties.
-A function which is at least $C^{2}$ is (strictly) convex if its hessian is positive (definite) semi-definite everywhere.
-If a Riemannian manifold admits a strictly convex function, then it has no compact immersed minimal submanifolds [20, §9.2 C].
+On simply connected Riemannian manifolds of nonpositive sectional curvature, called Hadamard manifolds, the square of the distance function from a fixed but arbitrary point and the Busemann functions are examples of convex functions which play a crucial role in the geometry of these manifolds.
+While the former is a smooth strictly convex function, the latter is merely a $C^{2}$-convex function [Ebe96, §1.6.6].
+In the nonnegative sectional curvature situation for instance, a key idea in the fundamental work [CG72] is the construction of suitable continuous convex functions related to Busemann functions and having appropriate additional properties.
+A function is (strictly) convex if its Hessian is positive (definite) semidefinite everywhere.
+If a Riemannian manifold admits a strictly convex function, then it has no compact immersed minimal submanifolds.
 In particular, it has no closed geodesics.
-Such a manifold also allows for a generalization of Cartan’s theorem [29] when this strictly convex function has a minimum: every compact subgroup of the isometry group has a common fixed point.
+When the strictly convex function has a minimum, such a manifold also allows for a generalisation of Cartan’s theorem [Yam82]: every compact subgroup of the isometry group has a common fixed point.
 The existence of strict convex functions also has consequences for the $\mathrm{L^{2}}$-spectrum of the Laplacian.
-The study of the spectrum of the Laplacian on complete Riemannian manifolds has a long, rich history and there are very many results, often with various additional constraints placed on the structure of the manifold and on sectional curvature.
-For a succint recent overview of some of these aspects, we refer to [9].
-We content ourselves by mentioning a few results which use convexity in an essential manner to deduce information on the spectrum to put our work in context, with no pretensions to completeness.
-The absolutely continuous part of the spectrum of the Laplacian on complete Riemannian manifolds supporting a sufficiently regular strictly convex function is nontrivial [28, Theorem 2] under certain boundedness conditions on the gradient and the Laplacian of the said function.
-On complete simply connected manifolds with nonpositive sectional curvature, [28, Theorem 1] exhibited a strictly convex function related to the square of the distance function satisfying the estimates required to ensure the nontriviality of the absolutely continuous spectrum.
-Moreover, it was also shown that in such a setting there are no eigenvalues in the spectrum.
-In the case of nonnegative sectional curvature, explicit convex functions have been used to show that there are no eigenfunctions under the additional assumptions on the codimension of the soul of the manifold being $1$ or having a quadratic decay of radial curvatures when the soul is just a point [14, Theorems 4.2, 4.3].
-One of the key results of this article is the construction of a $C^{2}$-strictly convex function with bounded gradient on complete simply connected Riemannian manifolds without focal points having negative Ricci curvature (Theorem 3.1).
-A distinguishing feature here is that we do not make any assumptions on the sign of the sectional curvature.
-It is well known [18] that there are both compact and noncompact manifolds without focal points whose sectional curvatures change sign.
-Since the Ricci curvature is an average of sectional curvatures, the sign of the Ricci curvature is a weaker notion when compared to the sign of the sectional curvature.
-That this is true especially for negative Ricci curvature can be seen from the remarkable result [23, Theorem A] that any smooth manifold of dimension at least three admits a complete Riemannian metric with negative Ricci curvature.
-Interesting classes of manifolds satisfying our assumptions include Riemannian symmetric spaces of noncompact type as well as simply connected noncompact nonflat harmonic manifolds without focal points.
-Harmonic manifolds are Einstein [21] i.e, $Ric_{g}=\kappa g$, and have horospheres, which are level sets of Busemann functions, of constant mean curvature $h\geq 0$.
+The study of the spectrum of the Laplacian on complete Riemannian manifolds has a long, rich history and there are very many results, often with various additional constraints placed on the structure of the manifold and on its curvature.
+For instance, in the case of manifolds of nonpositive curvature the essential spectrum was shown to be an infinite interval depending on the dimension and sectional curvatures at infinity in [Don81, Theorem 6.3] using the distance function.
+One can then ask interesting questions about the finer structure of the spectrum e.g., whether there are eigenvalues in the essential spectrum.
+For a succinct recent overview of some of these aspects, we refer to [Don10].
+We now content ourselves by mentioning a few results which use convexity in an essential manner to deduce refined information on the nature of the spectrum and to put our work in context, with no pretensions to completeness.
+In the case of nonnegative sectional curvature, explicit convex functions have also been used to show that there are no $\mathrm{L}^{2}$-eigenfunctions under the additional assumptions on the codimension of the soul of the manifold being $1$ or having a quadratic decay of radial curvatures when the soul is just a point [EF92, Theorems 4.2, 4.3].
+On Hadamard manifolds, [Xav88, Theorem 1] exhibited a strictly convex function related to the square of the distance function satisfying the estimates required to ensure the nontriviality of the absolutely continuous spectrum.
+However, this is not sufficient to exclude the existence either of eigenvalues or of the singularly continuous spectrum.
+In this article, we consider convex functions on complete Riemannian manifolds $M$ whose Jacobi fields satisfy certain growth properties, while making no assumptions on sectional curvatures.
+A simply connected Riemannian manifold is said to have no focal points if the norm of any nontrivial Jacobi field which vanishes at the initial point of a geodesic ray is strictly increasing.
+Since the exponential map at any point on such a manifold is a diffeomorphism, i.e.
+every point is a pole, they are necessarily noncompact.
+This class contains all Hadamard manifolds; however, it is well known [Gul75] that there are both compact and noncompact manifolds without focal points whose sectional curvatures change sign.
+While several aspects of the geometry of these manifolds have been well-studied, from understanding the dynamics of the geodesic flow [Ebe72], [O’S74], [Esc77]) to recent work on X-ray transforms on such manifolds with strictly convex boundary [PSUZ19] and the solution to the Dirichlet problem at infinity [LZ25], some aspects such as the spectrum of the Laplacian have not been analysed as much.
+We study some of these using the convexity of two kinds of geometrically defined functions on them.
+One of them is the square of the distance function from an arbitrary but fixed point.
+This is a smooth strictly convex function [Ebe72, Lemma 1] whose sub-level sets are compact.
+It is suitably modified to ensure a bounded gradient, an aspect that is pivotal in many applications.
+The other is the Busemann function associated to a unit vector in the tangent bundle of $M$.
+These are $C^{2}$ convex functions [Esc77, Theorem 2] but are not strictly convex.
+We remedy the situation by giving a new construction of a strictly convex function from Busemann functions.
+This leads to some new spectral theoretic results.
+A careful analysis of Jacobi fields and Jacobi tensors is central to these considerations.
+We now proceed to discuss our results in some detail.
+Our first application using the strict convexity of the suitably modified distance function concerns the isoperimetric profile function $\mathcal{I}_{M}$ defined on the manifold $M$.
+This function maps every positive volume to the infimum of the perimeters of sets in $M$ having that volume.
+The isoperimetric profile is shown [Rit17, Theorem 3.1] to be continuous and nondecreasing for complete manifolds which possess smooth strictly convex Lipschitz exhaustion functions and is therefore differentiable almost everywhere on such manifolds.
+Examples are discussed under sectional curvature conditions there, namely, Hadamard manifolds and manifolds with strictly positive sectional curvatures.
+In contrast, we observe in Proposition 3.2 that for simply connected Riemannian manifolds without focal points the map $\mathcal{I}_{M}$ is a nondecreasing continuous function, thus giving a class of examples without constraints on sectional curvatures.
+Strictly convex functions on Kähler manifolds can be seen to be strictly plurisubharmonic functions.
+So, manifolds without focal points which are also Kähler, admit strictly plurisubharmonic functions obtained, for instance, via the square of the distance function.
+From this, we observe in Proposition 3.4 that such manifolds are Stein using Grauert’s proof of Levi’s problem [GW77, §4].
+By composing with a suitable strictly convex function on $\mathbb{R}$, we get a strictly convex exhaustion function with bounded gradient.
+This is applied in Proposition 3.5 to obtain a lower bound for the growth of the volume of metric balls following [NT03, Lemma 4.2 (i)].
+The growth of the volume of the metric balls has been usually studied with lower bounds on Ricci curvature and upper bounds on sectional curvatures.
+It is an open problem whether a simply connected harmonic manifold which admits a Kähler structure is a symmetric space.
+Indeed in the Damek-Ricci setting this is known [BTV95, §4.1.12].
+Therefore, at least in certain special cases such as harmonic manifolds without focal points, this volume growth is in line with the above mentioned rigidity by Bishop-Günther’s comparison theorem [Cha06, Theorem III.4.2].
+We now proceed to describe our first spectral-theoretic result and note that the smooth strictly convex function with bounded gradient discussed in the previous paragraph is used here as well.
+A key idea is due to [Xav88] showing that on complete Riemannian manifolds supporting a sufficiently regular strictly convex function, Kato’s theory of $H$-smooth operators can be applied to understand the spectrum of the Laplacian.
+For a detailed discussion on $H$-smoothness we refer to the comprehensive survey [Sim19, §4].
+In the case of Hadamard manifolds, this was done in [Xav88, Theorem 1] under suitable bounds on the Laplacian and its derivatives to conclude that the absolutely continuous part of the spectrum was nontrivial.
+In fact, it contains an interval that depends explicitly on these bounds.
+In the present work, we extend the implementation of $H$-smoothing to simply connected Riemannian manifolds without focal points in Theorem 3.8.
+Assuming only the nonpositivity of radial curvatures with respect to a point, a condition that is markedly weaker than global nonpositivity of sectional curvature, and bounds on Laplacian and the bilaplacian, we show that the absolutely continuous spectrum is an infinite interval which depends on the bounds.
+This is carried out by applying Lemma 3.6 where we obtain the required quantitative estimates on the Hessian of distance function in the usual fashion, while using a function that is defined when the manifold necessarily has no focal points.
+For context, we mention some of the work that derive information about the spectrum from the behaviour of the radial curvatures.
+Absence of point spectrum has been shown for manifolds with a pole satisfying quadratic decay of radial curvatures [Don10, Theorem 1.5].
+Under linear decay of radial curvatures it was proved in [Kum10] that there are no eigenvalues embedded in the essential spectrum.
+If the radial curvature is assumed to approach a strictly negative constant near infinity, then [JL20] investigated the nature of the singular spectrum.
+In view of Theorem 3.8, we can ask for more refined properties of the spectrum such as identifying whether the point spectrum or the singularly continuous spectrum are nontrivial.
+A sufficient condition for the absence of point spectrum and the singularly continuous spectrum is given by the existence of a strictly convex function whose bilaplacian is nonpositive [Xav88, Corollary to Theorem 2].
+As noted there, this seemingly unnatural additional condition is, in fact, necessary.
+Indeed, if this condition on the bilaplacian fails, Hadamard manifolds with eigenvalues in the spectrum are constructed.
+This sufficiency condition seems to us not to be approachable by the use of distance function.
+To overcome this, we now turn our attention on the other natural geometrically defined convex function on simply connected Riemmanian manifolds without focal points, the Busemann function.
+Note that the Busemann function is convex but not strictly convex due to the vanishing of Hessian in the direction of its gradient.
+We use the geometry at infinity to construct a strictly convex function in Theorem 4.1 by averaging over all Busemann functions.
+This was motivated by the idea outlined in the context of symmetric spaces of noncompact type in [DX06, p-657], and is possibly of independent interest.
+For this, we make use of Jacobi tensors introduced in [Esc77] and its relation to the Hessian of Busemann functions for manifolds of continuous asymptote.
+In particular, for manifolds without focal points these Hessians can be seen to be positive semidefinite.
+The Hessian of the convex function which we construct by averaging over Buseman functions becomes positive definite under the assumption of negative Ricci curvature, yielding Theorem 4.1.
+We now examine the hypotheses of Theorem 4.1.
+In this class of manifolds, having nonnegative Ricci curvature at even one point, forces the manifold to be flat [MZ00, Corollary 1].
+Being an average of sectional curvatures, the sign of the Ricci curvature is a milder hypothesis.
+That this is true especially for negative Ricci curvature can be seen from the remarkable result [Loh94, Theorem A] that any smooth manifold of dimension at least three admits a complete Riemannian metric with negative Ricci curvature.
+Well-known classes of manifolds satisfying our assumptions include Riemannian symmetric spaces of noncompact type as well as simply connected noncompact (nonflat) harmonic manifolds without focal points.
+Harmonic manifolds are Einstein [Kni16] i.e, $Ric_{g}=\kappa g$, and have horospheres, which are level sets of Busemann functions, of constant mean curvature $h\geq 0$.
 If the Ricci curvature $\kappa>0$, then the harmonic manifold is necessarily a compact rank one symmetric space and if $\kappa=0$, then it is necessarily Euclidean.
 All other simply connected harmonic manifolds have negative Ricci curvature.
-We remark here that $h>0$ for such manifolds, since a harmonic manifold is Euclidean when $h=0$ [21, §5, Theorem 19].
-For the construction of the strictly convex function we make use of Jacobi tensors introduced in [13] and its relation to the Hessian of Busemann functions for manifolds of continuous asymptote.
-In particular, for manifolds without focal points these Hessians can be seen to be nonnegative.
-So each Busemann function is convex but not strictly convex since its Hessian vanishes in the direction of the gradient.
-Motivated by the work done in the context of symmetric spaces of noncompact type in [10], we then average the Busemann functions over the boundary of the geodesic compactification for manifolds without focal points.
-Here we use [16] for the geometry at infinity of such manifolds.
-Under the assumption of negative Ricci curvature, a strictly convex $C^{2}$-function with bounded gradient is then obtained.
-As an immediate consequence, we have that the manifolds under consideration admit smooth strictly convex functions which also has bounded gradient.
-This follows from the approximation result of Green and Wu [17, Theorem 2(a)].
-It is known [26] that the smooth approximation result fails for functions which are merely convex.
-Our next application is in the context of manifolds which are, in addition, Kähler.
-A complex manifold with a complex structure $J$ and a compatible Riemannian metric $g$ is called Kähler if the associated alternating two-form defined by $\omega(X,Y):=g(JX,Y)$ is symplectic.
-Given any two of these compatible structures the third one can be recovered.
-${\mathbb{C}}^{n}$ with the standard metric, bounded domains in ${\mathbb{C}}^{n}$ with the Bergmann metric, Hermitian symmetric spaces with the invariant metric, complex projective spaces with the Fubini-Study metric are some illustrative examples of Kähler manifolds.
-$C^{2}$-strictly convex functions on Kähler manifolds can be seen to be strictly plurisubharmonic functions.
-In our context, therefore, we get a smooth strictly plurisubharmonic function with bounded gradient.
-From this we derive a lower bound for the growth of the volume of metric balls on such manifolds following [24, Lemma 4.2 (i)].
-The growth of the volume of the metric balls has been usually studied with lower bounds on Ricci curvature and upper bounds on sectional curvatures [7, §III.4].
-As principal applications of Theorem 3.1, we describe the spectrum of the Laplacian for certain classes of manifolds having horospheres of constant mean curvature $h$.
-Complete simply connected Riemannian manifolds without conjugate points such that horospheres are of constant mean curvature $h\geq 0$ are known as asymptotically harmonic manifolds [21].
-On manifolds as in Theorem 3.1, which are in addition asymptotically harmonic, we observe that the strictly convex function that we construct also has bounded Laplacian and proceed to show in Theorem 4.3 that the spectrum is purely absolutely continuous and is contained in $[h^{2}/4,\infty)$.
-The condition on the bilaplacian required as in [28, Theorem 2] follows automatically since the horospheres have constant mean curvature.
-In fact, along the way we prove that equality holds in Cheeger’s inequality [6, §4.3, Theorem 3] for asymptotically harmonic manifolds and that Cheeger’s constant is equal to $h$ in Proposition 4.1.
-Indeed, the proof shows that it is sufficient for horospheres to have constant mean curvature in one direction.
-Further, we prove in Theorem 4.7 that the spectrum is purely absolutely continuous for simply connected noncompact nonflat harmonic manifolds without focal points and is equal to $[h^{2}/4,\infty)$.
-These are harmonic manifolds of purely exponential volume growth [21, Theorems 32, 45].
-Rank one symmetric spaces of noncompact type and other Damek-Ricci spaces are some examples of this class of manifolds.
-Since harmonic manifolds are asymptotically harmonic, that the spectrum is purely absolutely continuous and that it is contained in $[h^{2}/4,\infty)$ follow from Theorem 4.3.
-To prove that the spectrum is equal to $[h^{2}/4,\infty)$, we make use of the Fourier transform developed in this context in [3].
-More precisely, using the unitarity of the spherical Fourier transform [3, Theorem 4.7], we can show in the usual way that the multiplication operator $M_{(\lambda^{2}+\frac{h^{2}}{4})}$ is unitarily equivalent to the action of the Laplacian on radial functions to conclude that $[h^{2}/4,\infty)$ is contained in the spectrum.
-Higher rank symmetric spaces of noncompact type are not asymptotically harmonic since the horospheres along each direction have constant mean curvature which depend on that direction.
-Nonetheless, we are able to obtain (Theorem 4.8) that the spectrum of the Laplacian is purely absolutely continuous and is contained in $[\left\lVert\rho^{2}\right\rVert,\infty)$ by making use of the explicit expression for the mean curvature of horospheres and the strictly convex function that we construct in Theorem 3.1.
-To show that the spectrum is equal to $[\left\lVert\rho^{2}\right\rVert,\infty)$, we use the unitarity of just the spherical Fourier transform as in the case of harmonic manifolds.
-Note that in contrast to the case of harmonic manifolds where the unitarity of the Fourier transform is not known, it is classical that the Fourier transform in the case of symmetric spaces is unitary.
-As a consequence, the spectrum can also be completely determined for symmetric spaces via the unitary equivalence between the Laplacian and the multiplication operator $M_{(\left\lVert\lambda\right\rVert^{2}+\left\lVert\rho\right\rVert^{2})}$ given by the Fourier transform.
-It is worth remarking that in showing that the spectrum of the Laplacian on symmetric spaces of noncompact type is purely absolutely continuous, essential use is made only of the fact that they are manifolds without focal points and have negative Ricci curvature.
+Rank one symmetric spaces of noncompact type and other Damek-Ricci spaces are examples of this class of manifolds.
+We remark here that $h>0$ for such manifolds, since a harmonic manifold is Euclidean when $h=0$ [Kni16, §5, Theorem 19].
+Harmonic manifolds without focal points are of purely exponential volume growth, and these have been well studied in the recent past (e.g.[Kni16, BKP21, Bra24]).
+As principal applications of Theorem 4.1, we fully describe the nature of the spectrum $\sigma(\Delta)$ of the Laplacian $\Delta$ for interesting classes of manifolds having horospheres of constant mean curvature $h$.
+Simply connected Riemannian manifolds without conjugate points such that horospheres are of constant mean curvature $h\geq 0$ are known as asymptotically harmonic manifolds [Kni16].
+For all asymptotically harmonic manifolds, we show that equality holds in Cheeger’s inequality and that the Cheeger’s constant is equal to $h$ (see [KP15, §5] for related results).
+We mention in this context that for asymptotically harmonic Hadamard manifolds, the essential spectrum has been determined in [CS14, Theorem 4.4].
+But it did not address finer questions such as existence of embedded eigenvalues or the singularly continuous spectrum.
+It was proved in [BMP23] that for asymptotically harmonic Hadamard manifolds with strictly negative sectional curvature (and so necessarily without focal points) the spectrum is absolutely continuous.
+On simply connected manifolds without focal points and negative Ricci curvature, as in Theorem 4.1, which are asymptotically harmonic, we show in Theorem 4.3 that the spectrum is purely absolutely continuous and is contained in $h^{2}/4,\infty)$.
+Now since harmonic manifolds are asymptotically harmonic, the above result also holds in this case (Corollary 4.7).
+We give an elegant proof of the equality $\sigma(\Delta)=[h^{2}/4,\infty)$, making use of the Fourier transform developed recently in the context of harmonic manifolds without focal points in [BKP21].
+It is not known whether this Fourier transform is unitary on the space of all $\mathrm{L}^{2}$-functions.
+However, the Fourier transform is unitary on radial $\mathrm{L}^{2}$-functions [BKP21, Theorem 4.7] with the Plancherel density given by the $\mathrm{c}$-function, and intertwines the radial Laplacian and the multiplication operator $M_{(\lambda^{2}+\frac{h^{2}}{4})}$.
+Using certain estimates on the $\mathrm{c}$-function, we show that the essential range of the multiplication operator $M_{(\lambda^{2}+\frac{h^{2}}{4})}$ is equal to $h^{2}/4,\infty)$, enabling us to conclude the desired equality for the spectrum.
+Although symmetric spaces of higher rank are not asymptotically harmonic, our geometric methods allow us to recover the absolute continuity of the spectrum.
+It is, of course, classical that the unitarity of the Fourier transform in the case of symmetric spaces of noncompact type completely determines the spectrum of the Laplacian to be absolutely continuous and equal to $\left\lVert\rho^{2}\right\rVert,\infty)$.
 2. Preliminaries
-Let $M$ be a complete Riemannian manifold and let $\gamma$ be a geodesic in $M$. A vector field $Y$ on $\gamma$ is called a Jacobi field if it satisfies the Jacobi equation
-$Y^{{}^{\prime\prime}}+R(\gamma^{{}^{\prime}},Y)\gamma^{{}^{\prime}}=0.$
+In this section, we give a brief summary of the basic geometric setup needed in the sequel.
+Let $M$ be a complete Riemannian manifold and let $\gamma$ be a geodesic in $M$.
+A vector field $Y$ along $\gamma$ is called a Jacobi field if it satisfies the Jacobi equation
+(1) $Y^{{}^{\prime\prime}}+R(\gamma^{{}^{\prime}},Y)\gamma^{{}^{\prime}}=0.$
 Here $\gamma^{{}^{\prime}}$ is the tangent vector field of $\gamma$, $R$ is the Riemann curvature tensor on $M$ and $Y^{{}^{\prime}}$ denotes the covariant derivative of $Y$ with respect to $\gamma^{{}^{\prime}}$ along $\gamma$.
-Two points $p$ and $q$ on $\gamma$ are said to be conjugate along $\gamma$ if there exists a nonzero Jacobi field which vanishes both at $p$ and at $q$.
+The prototypical example of a Jacobi field is the variation vector field of a geodesic variation.
+Two points on $\gamma$ are said to be conjugate along $\gamma$ if there exists a nonzero Jacobi field which vanishes at both the points.
 $M$ is called a manifold without conjugate points if for every geodesic $\gamma$ in $M$, no two points on $\gamma$ are conjugate along $\gamma$ or equivalently, if the exponential map is non-singular at every point in $M$.
 In fact, $M$ has no conjugate points if and only if $\exp_{p}:T_{p}M\to M$ is a covering map for each $p$ in $M$.
-So if $M$ is also simply connected, then $\exp_{p}$ is a diffeomorphism for all $p\in M$ and hence $M$ is necessarily noncompact.
+So if $M$ is also simply connected, then $\exp_{p}$ is a diffeomorphism for all $p\in M$ [O’S74, Theorem 1] and hence $M$ is necessarily noncompact.
 In this case, any two distinct points of $M$ are joined by a unique geodesic.
-A general reference for the basic Riemannian geometry used in this article is [7].
-We now give a brief overview of Jacobi tensors and refer to [13] for a full account.
-Let $M$ be a complete simply connected Riemannian manifold without conjugate points and let ($SM$) $TM$ denote its (unit) tangent bundle.
+A simply connected Riemannian manifold $M$ is said to have no focal points if for every nontrivial Jacobi field $Y$ satisfying $Y(0)=0$, the map $t\mapsto\left\lVert Y(t)\right\rVert$ is strictly increasing for $t>0$ [O’S74, Proposition 4].
+In particular, any non-trivial Jacobi field cannot vanish more than once and hence every manifold without focal points is also a manifold without conjugate points.
+Note that all Hadamard manifolds are manifolds without focal points.
+For some applications, it is not just convenient but even important to consider the full vector space of normal Jacobi fields along a unit speed geodesic together rather than consider them one at a time.
+A Jacobi tensor is the tool which does this.
+We now give a brief overview of Jacobi tensors and refer to [Esc77] for a full account.
+Let $M$ be a simply connected Riemannian manifold without conjugate points and let ($SM$) TMTM denote its (unit) tangent bundle.
 For $v\in SM$, let $\gamma_{v}$ denote the unit speed geodesic determined by $v$.
@@
 A Jacobi tensor $Y_{v}$ is a normal $(1,1)$ - tensor field along a geodesic $\gamma_{v}$ with transversal derivative (i.e.
-there exists some $t$ such that $ker~{}Y_{v}(t)\cap ker~{}Y_{v}^{{}^{\prime}}(t)=0$) satisfying the Jacobi equation
-(2.1) $Y_{v}^{{}^{\prime\prime}}(t)+R(t)Y_{v}(t)=0$
+there exists some $t$ such that $ker\penalty 10000\ Y_{v}(t)\cap ker\penalty 10000\ Y_{v}^{{}^{\prime}}(t)=0$) satisfying the Jacobi equation
+(2) $Y_{v}^{{}^{\prime\prime}}(t)+R(t)Y_{v}(t)=0$
 Here $R(t)$ is the symmetric $(1,1)$-tensor along $\gamma_{v}$ induced by the curvature tensor $R$ and defined by
-$R(t)w:=R(w,\gamma_{v}^{{}^{\prime}}(t))\gamma_{v}^{{}^{\prime}}(t)$
+$R(t)w:=R(\gamma_{v}^{{}^{\prime}}(t),w)\gamma_{v}^{{}^{\prime}}(t)$
 and $Y_{v}^{{}^{\prime}}$ is the covariant derivative of the tensor $Y_{v}$ along $\gamma_{v}$ with respect to the tangent vector field $\gamma_{v}^{{}^{\prime}}$.
-For each $s>0$ and $v\in SM$, let $D_{s,v}$ be a solution to (2.1) with the boundary conditions $D_{s,v}(0)=\mathrm{Id}$ and $D_{s,v}(s)=0$.
+For each $s>0$ and $v\in SM$, let $D_{s,v}$ be a solution to the Jacobi equation (2) with the boundary conditions $D_{s,v}(0)=\mathrm{Id}$ and $D_{s,v}(s)=0$.
 Then $D_{v}:=\lim\limits_{s\to\infty}D_{s,v}$ exists and is referred to as the stable Jacobi tensor $D_{v}$ satisfying $D_{v}(0)=\mathrm{Id}$.
-A complete simply connected Riemannian manifold without conjugate points is said to be manifold of continuous asymptote if $v\to D_{v}^{{}^{\prime}}(0)$ is a continuous section of the vector bundle $\{(v,\phi):v\in SM,\phi\in\mathrm{End}(N(\gamma_{v})(0))\}$ over $SM$.
+A simply connected Riemannian manifold without conjugate points is said to be a manifold of continuous asymptote if $v\to D_{v}^{{}^{\prime}}(0)$ is a continuous section of the vector bundle $\{(v,\phi):v\in SM,\phi\in\mathrm{End}(N(\gamma_{v})(0))\}$ over $SM$.
+As stated explicitly in [Got79, p-596] simply connected manifolds without focal points fall in this class.
+For such manifolds, [KP13, §11] gives a compactification known as the Busemann compactification.
+Let $o\in M$ be an arbitrary but fixed point.
+Then the Busemann boundary $\partial^{o}M$ is homeomorphic to $S_{p}M$ for any $p\in M$ [KP13, Proposition 11.4 (2)].
+Convex functions being the theme of this article, the two natural geometrically defined functions which under suitable constraints on curvature are convex are the distance function and the Busemann function defined below.
 Given $t\in[0,\infty)$ and $v\in SM$, let $b_{t,v}:M\to\mathbb{R}$ denote the function defined by
@@
 The function $b_{v}$ is called the Busemann function associated to $v\in SM$.
-Recall that the level sets of Busemann functions are called horospheres.
+The level sets of Busemann functions are called horospheres.
 Horospheres are closed and hence are complete hypersurfaces with the induced metric.
-Recall that for Hadamard manifolds the Busemann functions are $C^{2}$ [12, §1.6.6].
-Eschenburg [13] was able to relax the curvature restriction by making use of convergence properties of Jacobi tensors, and generalized the $C^{2}$-regularity of horospheres to manifolds of continuous asymptote.
-For such manifolds, [22, §11] gives a compactification known as the Busemann compactification.
-Let $o\in M$ be an arbitrary but fixed point.
-Then the Busemann boundary $\partial^{o}M$ is homeomorphic to $S_{p}M$ for any $p\in M$ [22, Proposition 11.4 (2)].
-We see that for each $v\in SM$, the Hessian of the Busemann function $b_{v}$ and the stable Jacobi tensor $D_{v}$ are related as follows.
+It is classical that [Ebe96, §1.6.6] for Hadamard manifolds the Busemann functions are $C^{2}$ convex functions.
+Eschenburg [Esc77] was able to relax the curvature restriction by making use of convergence properties of Jacobi tensors, and generalized the $C^{2}$ - regularity of horospheres to manifolds of continuous asymptote.
+In order to obtain this, the following relation between the Hessian of the Busemann function $b_{v}$ and the stable Jacobi tensor $D_{v}$ for each $v\in SM$ is used.
+We first make a small observation.
 The orthogonal projection $x^{\perp}(p,v)$ of $x\in T_{p}M$ onto $(\nabla b_{v}(p))^{\perp}$ will be denoted by $x^{\perp}$ for notational convenience.
 Lemma 2.1.
-For $p\in M$ and $x\in S_{p}M$, we have ${\mathrm{H}}^{b_{v}}(p)(x,x)={\mathrm{H}}^{b_{v}}(p)(x^{\perp},x^{\perp})$.
+For $p\in M$ and $x\in S_{p}M$, we have $\mathrm{Hess}\penalty 10000\ b_{v}(p)(x,x)=\mathrm{Hess}\penalty 10000\ b_{v}(p)(x^{\perp},x^{\perp})$.
 Proof.
 As $\nabla b_{v}$ is a $C^{1}$ vector field with $\left\lVert\nabla b_{v}\right\rVert=1$, we have that $\nabla_{\nabla b_{v}}\nabla b_{v}=0$.
-Indeed ${\langle\nabla_{X}\nabla b_{v}},{Y\rangle}={\mathrm{H}}^{b_{v}}(X,Y)$ is symmetric in $X,Y$ and hence for every vector field $X$
+Indeed ${\langle\nabla_{X}\nabla b_{v}},{Y\rangle}=\mathrm{Hess}\penalty 10000\ b_{v}(X,Y)$ is symmetric in $X,Y$ and hence for every vector field $X$
 ${\langle\nabla_{\nabla b_{v}}\nabla b_{v}},{X\rangle}$ $={\langle\nabla_{X}\nabla b_{v}},{\nabla b_{v}\rangle}$
@@
 Then, the orthogonal decomposition $T_{p}M={\mathbb{R}}\nabla b_{v}\oplus(\nabla b_{v})^{\perp}$ enables us to write $x={\langle x},{\nabla b_{v}(p)\rangle}_{T_{p}M}\nabla b_{v}(p)+x^{\perp}$.
-Using $\nabla_{\nabla b_{v}}\nabla b_{v}=0$ and the bilinearity of ${\mathrm{H}}^{b_{v}}$ we have
-${\mathrm{H}}^{b_{v}}(p)(x,x)={\mathrm{H}}^{b_{v}}(p)(x^{\perp},x^{\perp}).$
-Notice in the above computation that when $X=\nabla b_{v}$ we have ${\mathrm{H}}^{b_{v}}(\nabla b_{v},\nabla b_{v})=0$ and hence $b_{v}$ is not strictly convex for any $v\in SM$.
+Using $\nabla_{\nabla b_{v}}\nabla b_{v}=0$ and the bilinearity of $\mathrm{Hess}\penalty 10000\ b_{v}$ we have
+$\mathrm{Hess}\penalty 10000\ b_{v}(p)(x,x)=\mathrm{Hess}\penalty 10000\ b_{v}(p)(x^{\perp},x^{\perp}).$
+Notice in the above computation that when $X=\nabla b_{v}$ we have $\mathrm{Hess}\penalty 10000\ b_{v}(\nabla b_{v},\nabla b_{v})=0$ and hence $b_{v}$ is not strictly convex for any $v\in SM$.
 In fact, to the best of our knowledge, it is not known if Busemann functions are even convex for an arbitrary manifold of continuous asymptote.
-Using Lemma 2.1 and [13, Proposition 2, Theorem 1] we have the following relation between the stable Jacobi tensors and the Busemann functions in the case of manifolds of continuous asymptote: for $v\in SM$, $p\in M$, and $x\in T_{p}M$,
-(2.2) ${\mathrm{H}}^{b_{v}}(p)(x,x)=-{\langle D_{\nabla b_{v}(p)}^{{}^{\prime}}(0)x^{\perp}},{D_{\nabla b_{v}(p)}(0)x^{\perp}\rangle}$
-Therefore for asymptotically harmonic manifolds, by taking traces in (2.2), we have $\Delta b_{v}=h=trace(D_{v}^{{}^{\prime}}(0))$ for every $v\in SM$.
-3. A strictly convex function on a manifold without focal points
-A complete simply connected Riemannian manifold $M$ is said to have no focal points if for every nontrivial Jacobi field $Y$ satisfying $Y(0)=0$, the map $t\mapsto\left\lVert Y(t)\right\rVert$ is strictly increasing for $t>0$ [25, Proposition 4].
-In particular, any non-trivial Jacobi field cannot vanish more than once and hence every manifold without focal points is also a manifold without conjugate points.
-Note that $t\mapsto\left\lVert D_{v}(t)\right\rVert$ is monotonically decreasing for each $v\in SM$.
-This in particular shows that $M$ is of continuous asymptote [13, §5].
-Henceforth in this section we assume that $M$ is a $n$-dimensional complete simply connected Riemannian manifold without focal points.
-To begin with let us recall the geodesic compactifications of $M$ and its connections to the Riemannian exponential map and the Busemann functions.
-We refer to [16] for more details on what follows.
-Given $v\in SM$ there exists a unique geodesic $\gamma_{v}$ in $M$ satisfying $\gamma_{v}^{{}^{\prime}}(0)=v$.
-Two vectors $v,w\in SM$ are said to be asymptotic if the distance between $\gamma_{v}(t)$ and $\gamma_{w}(t)$ is bounded for all $t\geq 0$.
-Let $M(\infty)$ be the set of all classes of asymptotic vectors and let $\overline{M}=M\cup M(\infty)$.
-$\overline{M}$ has a canonical topology with the following property: for any $p\in M$, the exponential map ${\mathrm{Exp}}_{p}:T_{p}M\to M$ extends uniquely to a homeomorphism $\Phi_{p}:\overline{T_{p}M}\to\overline{M}$ where $\overline{T_{p}M}\backslash T_{p}M$ is homeomorphic with the unit sphere $S_{p}M$ in $T_{p}M$.
-Then for any $p\in M$, there exists a unique geodesic $\gamma_{{}_{\nabla b_{v}(p)}}$ starting at $p$ asymptotic to $\gamma_{v}$.
-From our discussions in §2, we see that the Busemann boundary $\partial^{o}M$ and the geodesic compactification $M(\infty)$ are homeomorphic.
-Further, for each $v\in SM$, the Busemann function $b_{v}$ is convex [13, Theorem 2] i.e., ${\mathrm{H}}^{b_{v}}$ is nonnegative definite for every $v\in SM$.
-Hence each horosphere is a convex submanifold in $M$.
-As remarked in the introduction, motivated by the construction on symmetric spaces of noncompact type of a strictly convex function, indicated in [10, penultimate paragraph p.
-657], we now construct a strictly convex function out of the Busemann functions, using the geometry at infinity for manifolds without focal points discussed above.
+Using Lemma 2.1 and [Esc77, Proposition 2, Theorem 1] we have the following relation between the stable Jacobi tensors and the Busemann functions in the case of manifolds of continuous asymptote: for $v\in SM$, $p\in M$, and $x\in T_{p}M$,
+(3) $\mathrm{Hess}\penalty 10000\ {b_{v}}(p)(x,x)=-{\langle D_{\nabla b_{v}(p)}^{{}^{\prime}}(0)x^{\perp}},{D_{\nabla b_{v}(p)}(0)x^{\perp}\rangle}$
+In what follows, we shall make use of the foregoing discussion for simply connected Riemannian manifolds without focal points and henceforth make this assumption.
+3. A construction using the distance function and its applications
+In this section, we consider the distance from a fixed point on a simply connected manifold $M$ without focal points, and the strictly convex functions derived from it, having additional properties required for our applications.
+For a fixed $p\in M$, the function $q\mapsto d^{2}(p,q)$ is smooth and strictly convex where $r(q)=d(p,q)$ is the distance function.
+This is proved by an analysis of Jacobi fields as follows.
+Note that the normalized distance-squared function $u(q)=\frac{1}{2}r^{2}(q)$ is smooth on $M$.
+The Hessian of $u$ is given by
+$\mathrm{Hess}\penalty 10000\ u=\mathrm{d}r^{2}+r\mathrm{Hess}\penalty 10000\ r$
+Given $v\in T_{p}M$, consider the Jacobi field along $t\mapsto\exp_{p}(tv)$ satisfying $Y(0)=0$. Then a routine calculation yields,
+(4) $\mathrm{Hess}\penalty 10000\ r(\exp_{p}(tv))(Y(t),Y(t))={\langle Y^{{}^{\prime}}(t)},{Y(t)\rangle}=\frac{1}{2}\frac{d}{ds}\Big|_{s=t}\left\lVert Y(s)\right\rVert^{2}$
+Since $M$ has no focal points, $d\exp_{p}$ is nonsingular at $tv$ for any $w\in T_{q}M$ there exists a Jacobi field $Y$ along $\gamma$ such that $Y(0)=0$ and $Y(t)=w$.
+Using [O’S74, Theorem 4] we see that $\mathrm{Hess}\penalty 10000\ r>0$ for $t>0$ and so $u$ is strictly convex on $M$ and we refer to [Ebe72] for more details.
+It then follows from [Yam82, Theorem A] that each compact subgroup of the isometry group of $M$ has a common fixed point.
+This generalizes the classical theorem of Cartan for Hadamard manifolds.
+Now we construct, in the usual way, a smooth strictly convex function on $M$ whose gradient is bounded by taking $f(q)=\sqrt{1+2u(q)}$; then $\left\lVert\nabla f\right\rVert\leq 1$ and
+$\mathrm{Hess}\penalty 10000\ f$ $=\frac{1}{(1+2u)^{\frac{3}{2}}}\mathrm{d}r^{2}+\sqrt{\frac{2u}{1+2u}}\mathrm{Hess}\penalty 10000\ r$
+is positive definite as $\mathrm{Hess}\penalty 10000\ r$ is strictly positive showing that $f$ is strictly convex.
+A real-valued function $\varphi$ on a manifold is called an exhaustion function if all its sub-level sets $\{\varphi\leq c\}$ are compact.
+Notice that the sub-level sets of $f$ are given by
+$f^{-1}(c)=\begin{cases}\emptyset,&\text{if}\penalty 10000\ c<1\\ \{p\},&\text{if}\penalty 10000\ c=1\\ B(p,c),&\text{if}\penalty 10000\ c>1\end{cases}$
+and hence are all compact. Here $B(p,c)$ denotes the closed ball of radius $c$ centred at $p$. Thus $f$ is also an exhaustion function on $M$. For convenience, we summarize this discussion as a lemma.
+Lemma 3.1.
+Let $M$ be a simply connected Riemannian manifold without focal points. Then it possesses a smooth strictly convex exhaustion function with bounded gradient.
+As a first application we have the following result on the isoperimetric profile function $\mathcal{I}_{M}$ on $M$.
+The isoperimetric profile of a Riemannian manifold is the function that assigns to a given positive volume the infimum of the perimeter of the sets of this volume.
+We remark here that regularity properties of this function are of interest and refer to the introduction in [Rit17] for a detailed discussion.
+The continuity and the nondecreasing nature of the isoperimetric profile is shown [Rit17, Theorem 3.1] for complete manifolds which possess smooth strictly convex Lipschitz exhaustion functions.
+It then follows from a theorem of Lebesgue that $\mathcal{I}_{M}$ is differentiable almost everywhere on such manifolds.
+As examples there, Hadamard manifolds and manifolds with strictly positive sectional curvatures are discussed.
+We give below a class of examples for which the continuity of $\mathcal{I}_{M}$ holds while making no assumptions on curvature and which properly contains Hadamard manifolds.
+As noted before this class contains plenty of examples whose sectional curvatures do not have a fixed sign [Gul75].
+Proposition 3.2.
+Simply connected Riemannian manifolds without focal points have a continuous nondecreasing isoperimetric profile function. In particular, they are differentiable almost everywhere.
+Proof.
+By Lemma 3.1, $M$ admits a smooth strictly convex Lipschitz exhaustion function. Invoking [Rit17, Theorem 3.1], we see that $\mathcal{I}_{M}$ is nondecreasing and continuous. ∎
+Regularity for $\mathcal{I}_{M}$ has many interesting consequences; for instance, we have the following result whose proof makes use of standard arguments.
+Proposition 3.3.
+Let $M$ be as in Proposition 3.2.
+Suppose $v_{0}$ is a differentiable point of $\mathcal{I}_{M}$.
+If $\Omega$ is an isoperimetric region in $M$ of volume $v_{0}$, then (the regular part of) its boundary $\partial\Omega$ has constant mean curvature $H=\mathcal{I}_{M}^{{}^{\prime}}(v_{0})$.
+Proof.
+Since $\Omega$ is isoperimetric region the (regular part of) $\partial\Omega$ has constant mean curvature $H$.
+Let $\mu,\nu$ be the outward and inward normal variational vector fields on $\partial\Omega$ with compact support.
+Then the first variations of the volumes and the boundary areas are as follows
+$\frac{d}{dt}V(\Omega_{t})\Big|_{t=0}$ $=\int\limits_{\partial\Omega_{0}}|\mu|\mathrm{d}A,\quad\frac{d}{dt}V(\Omega_{t})\Big|_{t=0}=-\int\limits_{\partial\Omega_{0}}|\nu|\mathrm{d}A$
+$\frac{d}{dt}A(\partial\Omega_{t})\Big|_{t=0}$ $=H\int\limits_{\partial\Omega_{0}}|\mu|\mathrm{d}A,\quad\frac{d}{dt}A(\partial\Omega_{t})\Big|_{t=0}=-H\int\limits_{\partial\Omega_{0}}|\nu|\mathrm{d}A$
+Hence, using the differentiability assumption at $v_{0}$,
+$\lim\limits_{\delta v\to 0^{+}}\frac{\mathcal{I}_{M}(v_{0}+\delta v)-\mathcal{I}_{M}(v_{0})}{\delta v}\leq H,\quad\lim\limits_{\delta v\to 0^{-}}\frac{\mathcal{I}_{M}(v_{0}+\delta v)-\mathcal{I}_{M}(v_{0})}{\delta v}\geq H$
+yields $\mathcal{I}_{M}^{{}^{\prime}}(v_{0})=H$. ∎
+As a consequence, all isoperimetric regions in $M$ of the same volume $v_{0}$ must also have the same mean curvature if $v_{0}$ is a differentiable point of the isoperimetric profile function.
+In general, the existence of isoperimetric domains of a given volume in a complete noncompact Riemannian manifold is a difficult problem (see [Rit23]).
+We now discuss a couple of direct applications in the situation when $M$ is, in addition, Kähler.
+For general facts on Kähler manifolds we refer to [GW77, §7].
+On a Kähler manifold the complex analogue of convex functions is given by the notion of plurisubharmonic functions.
+These are defined using the complex Hessian called the Levi form.
+In local coordinates, the Levi form of a function $\tilde{f}$ is given by
+$L\tilde{f}=4\sum\limits_{i,j=1}^{n}\frac{\partial^{2}h}{\partial z_{i}\partial\overline{z}_{j}}\mathrm{d}z_{i}\otimes\mathrm{d}\overline{z}_{j}$
+where $(z_{1},\cdots,z_{n})$ is a local holomorphic coordinate system.
+Let $(x_{1},y_{1},\cdots,x_{n},y_{n})$ be the corresponding normal coordinates with $z_{k}=x_{k}+iy_{k}$, $k=1,\cdots,n$.
+A real-valued function on a Kähler manifold is said to be (strictly) plurisubharmonic, if the Levi form is positive (definite) semi-definite.
+Since $M$ is Kähler, the Levi form $L_{\tilde{f}}$ of $\tilde{f}$ is expressed in terms of Hessian of $\tilde{f}$ using the real coordinates as
+$L_{\tilde{f}}=\mathrm{Hess}(\tilde{f})\left(\frac{\partial}{\partial x_{1}},\frac{\partial}{\partial x_{1}}\right)+\mathrm{Hess}(\tilde{f})\left(\frac{\partial}{\partial y_{1}},\frac{\partial}{\partial y_{1}}\right).$
+Therefore if $\tilde{f}$ is a strictly convex $C^{2}$-function on a Kähler manifold, then it is strictly plurisubharmonic.
+Stein manifolds are complex manifolds which are biholomorphic to a closed complex submanifold of $\mathbb{C}^{n}$ for some $n\in\mathbb{N}$.
+This tells us that there are plenty of holomorphic functions on such manifolds.
+Stein manifolds are necessarily Kähler and a question of great interest then is to ask which Kähler manifolds are Stein.
+Typically, sufficient conditions are given by putting constraints on the sign of of sectional, holomorphic bisectional or Ricci curvatures[GW77, Theorem $1$].
+It is in this context that we observe the following sufficient condition where we make no assumptions on curvature.
+Proposition 3.4.
+Let $M$ be a simply connected Riemannian manifold without focal points which is also Kähler. Then $M$ is necessarily Stein.
+Proof.
+Lemma 3.1 gives a smooth strictly convex exhaustion function on $M$, while the discussion above on the Levi form shows that this function is a strictly plurisubharmonic exhaustion function on $M$.
+Then Grauert’s proof of Levi’s problem [GW77, §4] that a complex manifold is Stein if and only if it admits a strictly plurisubharmonic exhaustion function concludes the proof.
+∎
+As another application of Lemma 3.1, we have the following volume growth comparison for metric balls in Kähler manifolds without focal points using [NT03, Lemma 4.2 (i)].
+Proposition 3.5.
+Let $M$ be a simply connected Riemannian manifold without focal points which is also Kähler and of (complex) dimension $m$.
+Then $M$ admits a smooth strictly plurisubharmonic function with bounded gradient.
+Moreover, for each $p\in M$ there exists a constant $C_{2}>0$ such that,
+$V_{p}(r)\geq C_{2}r^{m},\quad r\geq 1$
+where $V_{p}(r)$ is the volume of a metric ball of radius $r$ about $p$ in $M$.
+As mentioned in the introduction, the question of whether a simply connected harmonic manifold that carries a Kähler structure must be a symmetric space is still unresolved.
+However, in the context of Damek–Ricci spaces this has been verified.
+Therefore, at least in certain special cases such as harmonic manifolds without focal points, the resulting volume growth reflects the same rigidity, in accordance with Bishop–Günther’s comparison theorem [Cha06, Theorem III.4.2].
+Our next application is spectral theoretic in nature, and so we begin with a sketch of the basic functional analytic setup concerning the spectrum of the Laplacian referring to [Bor20] for the details.
+Let $M$ be a complete Riemannian manifold.
+Then the (positive) Laplacian $\Delta$ is an essentially self-adjoint operator with the space of compactly supported smooth functions on $M$ as the domain.
+We also denote its unique self-adjoint extension, having the domain $\mathcal{D}(\Delta):=\{f\in L^{2}(M):\Delta f\in L^{2}(M)\}$, by $\Delta$.
+Using the functional calculus for $\Delta$ we have the family of spectral projection operators $\prod_{E}:=\chi_{E}(\Delta)$ for each Borel subset $E$ of $\mathbb{R}$} referred to as the resolution of $\Delta$.
+The support of these projections and the spectrum of $\Delta$ are related in the following way.
+We say that $\lambda\in{\mathbb{C}}$ is in the spectrum $\sigma(\Delta)$ of $\Delta$ if and only if $\prod_{(\lambda-\epsilon,\lambda+\epsilon)}\neq 0$ for all $\epsilon>0$.
+If $\prod_{\{\lambda\}}\neq 0$, then $\lambda$ is an eigenvalue of $\Delta$ and the range of $\prod_{\{\lambda\}}$ is the corresponding eigenspace.
+The spectral theorem gives a unitary equivalence between $\Delta$ and a multiplication operator on a suitable Hilbert space of the form $\mathrm{L}^{2}(X,\mu)$, where $\mu$ is a $\sigma$- finite Borel measure on a measure space $X$.
+By the Lebesgue decomposition, $\mu$ decomposes as $\mu=\mu_{pp}+\mu_{ac}+\mu_{sc}$ into pure point measure, absolutely continuous and singularly continuous measures with respect to the Lebesgue measure, and consequently we have that
+(5) $\mathrm{L}^{2}(X,\mu)=\mathrm{L}^{2}(X,\mu_{pp})\oplus\mathrm{L}^{2}(X,\mu_{ac})\oplus\mathrm{L}^{2}(X,\mu_{sc}).$
+This induces a decomposition $\mathrm{L}^{2}(M)=\mathcal{H}_{pp}\oplus\mathcal{H}_{ac}\oplus\mathcal{H}_{sc}$ from which we have the associated decomposition of the spectrum of $\Delta$ namely, $\sigma(\Delta)=\overline{\sigma_{pt}(\Delta)}\cup\sigma_{ac}(\Delta)\cup\sigma_{sc}(\Delta)$ where
+$\sigma_{ac}(\Delta)=\sigma(\Delta_{|_{\mathcal{H}_{ac}}}),\quad\sigma_{sc}(\Delta)=\sigma(\Delta_{|_{\mathcal{H}_{sc}}}),\quad\overline{\sigma_{pt}(\Delta)}=\sigma(\Delta_{|_{\mathcal{H}_{pp}}})$
+Here we take the closure for the set of all eigenvalues as it is not necessarily closed.
+We refer to $\sigma_{ac}(\Delta)$ and $\sigma_{sc}(\Delta)$ as the absolutely continuous and singularly continuous parts of $\sigma(\Delta)$ respectively and their union as the continuous part $\sigma_{cont}(\Delta)$ of $\sigma(\Delta)$.
+There is also a decomposition of $\sigma(\Delta)$ as a disjoint union:
+$\sigma(\Delta)=\sigma_{disc}(\Delta)\cup\sigma_{ess}(\Delta)$
+where $\sigma_{disc}(\Delta)$ and $\sigma_{ess}(\Delta)$ denote the discrete and essential parts of $\sigma(\Delta)$.
+The former consists of all eigenvalues of finite multiplicities and hence is contained in $\sigma_{pt}(\Delta)$.
+The latter is the complement of $\sigma_{disc}(\Delta)$ in $\sigma(\Delta)$ thereby we have that the continuous part of the spectrum, namely $\sigma_{ac}(\Delta)\cup\sigma_{sc}(\Delta)$ is contained in $\sigma_{ess}(\Delta)$.
+Note that it is possible that $\sigma_{ess}(\Delta)$ contains eigenvalues of infinite multiplicities usually referred to as the embedded eigenvalues.
+In order to analyse the nature of the spectrum of the Laplacian using convexity, we make use of [Xav88, Theorem 3].
+The idea there was to use Kato’s $H$-smooth operators to show that if $M$ admits a strictly convex function, then certain boundedness conditions on its gradient and Laplacian implies the non-triviality of the absolutely continuous part of the spectrum.
+This is then applied in the context of Hadamard manifolds with certain additional assumptions which are satisfied with lower bounds on sectional curvatures, to show the nontriviality of the absolutely continuous spectrum.
+It is this result that we now generalise to simply connected Riemannian manifolds without focal points having nonpositive radial curvature with respect to a single point.
+This is a much weaker condition than the nonpositivity of sectional curvature.
+To this end, we prove the following lemma.
+Lemma 3.6.
+Let $M$ be a simply connected Riemannian manifold without focal points.
+Suppose there exists a point $p$ in $M$ such that the radial curvatures with respect to $p$ are nonpositive.
+Then we have for any $q\neq p$ in $M$
+$r(q)\mathrm{Hess}\penalty 10000\ r(q)(v,v)\geq\left\lVert v\right\rVert^{2}\qquad\forall\penalty 10000\ v\in T_{q}M$
+Proof.
+Since $M$ has no focal points $\exp_{p}:T_{p}M\to M$ is a diffeomorphism.
+Hence for any $q\in M$, $q\neq p$, there exists a unique $t\in{\mathbb{R}}$ and a unit tangent vector $v\in T_{p}M$ such that $q=\exp_{p}(tv)$.
+Moreover, $\gamma(s)=\exp_{p}(sv)$ is a unit speed geodesic in $M$ joining $p$ and $q$ satisfying $\gamma(0)=p$ and $\gamma(t)=q$.
+Then $\gamma^{\prime}(s)=\nabla r(\gamma(s))$ and in particular, $\gamma^{\prime}(t)=\nabla r(q)$.
+Since $d\exp_{p}$ is nonsingular at $tv$, for any $w\in T_{q}M$ there exists a Jacobi field $Y$ along $\gamma$ such that $Y(0)=0$ and $Y(t)=w$.
+Notice that by our assumption on nonpositive radial curvature, we have ${\langle R(Y(t),\gamma^{{}^{\prime}}(t))\gamma^{{}^{\prime}}(t)},{Y(t)\rangle}\leq 0$ for any $t>0$ and $v\in S_{p}M$.
+We now recall the standard technique of Jacobi field comparison.
+Since the second derivative of the Jacobi field is related to the curvature tensors via the Jacobi equation, quantitative estimates are obtained by imposing conditions on curvatures.
+Note that
+$\lambda(s)=\frac{\left\lVert Y(s)\right\rVert^{2}}{{\langle Y^{{}^{\prime}}(s)},{Y(s)\rangle}},\penalty 10000\ s>0$
+is well defined since $M$ has no focal points. Using l’Hospital rule, we see that $\lambda(0)=0$ since $Y^{{}^{\prime}}(0)\neq 0$. Differentiating and using the Jacobi equation we have
+$\lambda^{{}^{\prime}}(s)$ $=\frac{2{\langle Y^{{}^{\prime}}(s)},{Y(s)\rangle}^{2}-\left(\left\lVert Y^{{}^{\prime}}(s)\right\rVert^{2}-{\langle R(\gamma^{{}^{\prime}}(s),Y(s))\gamma^{{}^{\prime}}(s)},{Y(s)\rangle}\right)\left\lVert Y(s)\right\rVert^{2}}{{\langle Y^{{}^{\prime}}(s)},{Y(s)\rangle}^{2}}$
+$\leq\frac{2{\langle Y^{{}^{\prime}}(s)},{Y(s)\rangle}^{2}-\left\lVert Y^{{}^{\prime}}(s)\right\rVert^{2}\left\lVert Y(s)\right\rVert^{2}}{{\langle Y^{{}^{\prime}}(s)},{Y(s)\rangle}^{2}}\quad\text{(nonpositivity of radial curvature)}$
+$\leq\frac{2{\langle Y^{{}^{\prime}}(s)},{Y(s)\rangle}^{2}-{\langle Y^{{}^{\prime}}(s)},{Y(s)\rangle}^{2}}{{\langle Y^{{}^{\prime}}(s)},{Y(s)\rangle}^{2}}=1\quad\text{(Cauchy-Schwarz inequality)}$
+Hence $\lambda(s)\leq s$ or equivalently,
+$s{\langle Y^{{}^{\prime}}(s)},{Y(s)\rangle}\geq\left\lVert Y(s)\right\rVert^{2},\quad s>0$
+The claim follows from (4). ∎
+We are now ready to prove the main spectral result of this section.
+The bounds on the Laplacian and the bilaplacian demanded below are satisfied if lower bounds on curvatures and derivatives of metric tensor are assumed.
+It should be kept in mind that there exist Hadamard manifolds with pure point spectrum if the curvatures become unbounded.
+Remark 3.7.
+Note that we adopt the convention of considering the Laplacian as a positive operator, in contrast to the convention used in [Xav88] for Hadamard manifolds.
+Consequently, certain inequalities in the subsequent theorem appear with opposite signs when compared to the corresponding proof in [Xav88].
 Theorem 3.8.
-Let $M$ be a complete non-compact simply connected Riemannian manifold with no focal points and negative Ricci curvature.
-Then $M$ admits a $C^{2}$-strictly convex function $F$ whose gradient is bounded.
+Let $M$ be a simply connected Riemannian manifold without focal points.
+Suppose there exists a point $p$ in $M$ such that the radial curvatures with respect to $p$ are nonpositive.
+If there exist constants $c_{1},c_{2}>0$ such that $f$ and the Laplacian $\Delta$ satisfies
+  1. (i)
+    $\Delta f\geq-c_{1}$
+  2. (ii)
+    $\Delta^{2}f\leq c_{2}f^{-3}$
+then the absolutely continuous spectrum of $\Delta$ is nontrivial.
+In fact, $(\alpha,\infty)$ is in the absolutely continuous part of the spectrum of $\Delta$ where $\alpha=\frac{3}{2}c_{1}+\frac{1}{4}c_{2}$
+Proof.
+Note that $\left\lVert\nabla f\right\rVert$ is bounded.
+Since $M$ has nonpositive radial curvature at $p$, from Lemma 3.6 we obtain that $\mathrm{Hess}\penalty 10000\ r^{2}(p)$ is the Riemannian metric at $p$ and for any $p\neq q\in M$,
+$\mathrm{Hess}\penalty 10000\ r^{2}(q)(w,w)={\langle\nabla r(q)},{w\rangle}^{2}+r(q)\mathrm{Hess}\penalty 10000\ r(q)(w,w)\geq\left\lVert w\right\rVert^{2},\quad\forall\penalty 10000\ w\in T_{q}M,\left\lVert w\right\rVert=1$
+From here, the arguments are standard, as in [Xav88, Theorem 1], for example. We write them down here for completeness. Recall that $f(q)=(1+r(q)^{2})^{1/2}$ and its Hessian can be computed to be
+$\mathrm{Hess}\penalty 10000\ f(q)(w,w)$ $=-\frac{r^{2}(q)}{(1+r^{2}(q))^{3/2}}{\langle w},{\nabla r(q)\rangle}^{2}+\frac{1}{2(1+r^{2}(q))^{1/2}}\mathrm{Hess}\penalty 10000\ r^{2}(q)(w,w)$
+$\geq\frac{1}{(1+r^{2}(q))^{3/2}}$
+Taking $h(q)=\frac{1}{(1+r^{2}(q))^{3/2}}$, we see that
+$\Delta h$ $=\frac{3}{f^{4}}\Delta f+\frac{12}{f^{5}}\left\lVert\nabla f\right\rVert^{2},$
+$\implies\Delta h$ $\geq\frac{3\Delta f}{f^{4}}\geq-\frac{3c_{1}}{2}$
+Thus for
+$\beta=\inf\frac{1}{4h}\{2\Delta h-\Delta^{2}f\}\geq-\frac{3c_{1}}{2}-\frac{c_{2}}{4}$
+and the claim now follows as in [Xav88, Theorem 3]. ∎
+Clearly, the above theorem generalizes [Xav88, Theorem 1] under much weaker curvature conditions.
+However, more refined questions such as identifying whether the point spectrum or the singularly continuous spectrum occur are not amenable to this approach via the use of the distance function.
+To overcome this, we now turn our attention on the other natural geometrically defined convex function, the Busemann function.
+4. A construction using Busemann functions and its applications
+In this section, we construct a smooth strictly convex function out of Busemann functions on simply connected manifolds without focal points, using the geometry at infinity and assuming, in addition, strict negativity of Ricci curvature.
+As remarked in the introduction, this is motivated by a similar construction on symmetric spaces of noncompact type outlined in [DX06, penultimate paragraph p.657657].
+The main purpose of this construction here is that it allows us to prove new results (Theorem 4.3 and Corollary 4.7) completely determining the nature of the spectrum for certain interesting classes of manifolds without focal points whose horospheres have constant mean curvature.
+[…]
+Theorem 4.1.
+Let $M$ be simply connected Riemannian manifold with no focal points and negative Ricci curvature.
+Then $F(p)=\displaystyle{\int\limits_{M(\infty)}b_{v}(p)\penalty 10000\ \mathrm{d}\eta(v)}$ is a strictly convex function whose gradient is bounded.
 Proof.
@@
 We identify $S_{o}M$, $M(\infty)$ and $\partial^{o}M$ via the respective homeomorphisms.
-Let $F(p)=\displaystyle{\int\limits_{M(\infty)}b_{v}(p)~{}\mathrm{d}\eta(v)}$.
-By dominated convergence theorem, the gradient of $F$ at $p$ is given by $\nabla F(p)=\displaystyle{\int\limits_{M(\infty)}\nabla b_{v}(p)~{}\mathrm{d}\eta(v)}$.
+By dominated convergence theorem, the gradient of $F$ at $p$ is given by $\nabla F(p)=\displaystyle{\int\limits_{M(\infty)}\nabla b_{v}(p)\penalty 10000\ \mathrm{d}\eta(v)}$.
 Clearly,
-$\left\lVert\nabla F(p)\right\rVert=\left\lVert\int\limits_{M(\infty)}\nabla b_{v}(p)~{}\mathrm{d}\eta(v)\right\rVert\leq\int\limits_{M(\infty)}\left\lVert\nabla b_{v}(p)\right\rVert~{}\mathrm{d}\eta(v)=1,\quad\text{as}~{}\left\lVert\nabla b_{v}\right\rVert=1$
-i.e., $F$ has bounded gradient. In particular, $F$ is Lipschitz with Lipschitz constant $1$. For $p\in M$ and $w\in S_{p}M$, using (2.2) and that ${\mathrm{H}}^{b_{v}}$ is nonnegative definite we have
-$\int\limits_{M(\infty)}{\mathrm{H}}^{b_{v}}(p)(w,w)~{}\mathrm{d}\eta(v)$ $=\left|\int\limits_{M(\infty)}{\langle D_{\nabla b_{v}(p)}^{{}^{\prime}}(0)w},{w\rangle}~{}\mathrm{d}\eta(v)\right|$
-$\leq\int\limits_{M(\infty)}|{\langle D_{\nabla b_{v}(p)}^{{}^{\prime}}(0)w},{w\rangle}|~{}\mathrm{d}\eta(v)$
-[…]
-$\leq\int\limits_{M(\infty)}\left\lVert D_{\nabla b_{v}(p)}^{{}^{\prime}}(0)\right\rVert~{}\mathrm{d}\eta(v)\quad(D_{\nabla b_{v}(p)}^{{}^{\prime}}(0)~{}\text{is a bounded linear map for}~{}v\in S_{p}M).$
-The restriction of the continuous map $w\mapsto\left\lVert D_{w}^{{}^{\prime}}(0)\right\rVert$ on $SM$ to $S_{p}M$ shows that there exists some $L>0$ such that $\left\lVert D_{w}^{{}^{\prime}}(0)\right\rVert\leq L$ for all $w\in S_{p}M$.
+$\left\lVert\nabla F(p)\right\rVert=\left\lVert\int\limits_{M(\infty)}\nabla b_{v}(p)\penalty 10000\ \mathrm{d}\eta(v)\right\rVert\leq\int\limits_{M(\infty)}\left\lVert\nabla b_{v}(p)\right\rVert\penalty 10000\ \mathrm{d}\eta(v)=1,\quad\text{as}\penalty 10000\ \left\lVert\nabla b_{v}\right\rVert=1$
+i.e., $F$ has bounded gradient.
+In particular, $F$ is Lipschitz with Lipschitz constant $1$.
+For $p\in M$ and $w\in S_{p}M$, using (3) and that $\mathrm{Hess}\penalty 10000\ b_{v}$ is positive semidefinite we have
+$\int\limits_{M(\infty)}\mathrm{Hess}\penalty 10000\ b_{v}(p)(w,w)\penalty 10000\ \mathrm{d}\eta(v)$ $=\left|\int\limits_{M(\infty)}{\langle D_{\nabla b_{v}(p)}^{{}^{\prime}}(0)w},{w\rangle}\penalty 10000\ \mathrm{d}\eta(v)\right|$
+$\leq\int\limits_{M(\infty)}|{\langle D_{\nabla b_{v}(p)}^{{}^{\prime}}(0)w},{w\rangle}|\penalty 10000\ \mathrm{d}\eta(v)$
+$\leq\int\limits_{M(\infty)}\left\lVert D_{\nabla b_{v}(p)}^{{}^{\prime}}(0)w\right\rVert\left\lVert w\right\rVert\penalty 10000\ \mathrm{d}\eta(v)\quad\text{(using Cauchy-Schwarz)}$
+$\leq\int\limits_{M(\infty)}\left\lVert D_{\nabla b_{v}(p)}^{{}^{\prime}}(0)\right\rVert\penalty 10000\ \mathrm{d}\eta(v)\quad(D_{\nabla b_{v}(p)}^{{}^{\prime}}(0)\penalty 10000\ \text{is a bounded linear map for}\penalty 10000\ v\in S_{p}M).$
+The restriction of the continuous map (§2) $w\mapsto\left\lVert D_{w}^{{}^{\prime}}(0)\right\rVert$ on $SM$ to $S_{p}M$ shows that there exists some $L>0$ such that $\left\lVert D_{w}^{{}^{\prime}}(0)\right\rVert\leq L$ for all $w\in S_{p}M$.
 In particular, $\left\lVert D_{\nabla b_{v}(p)}^{{}^{\prime}}(0)\right\rVert\leq L$ for all $v\in S_{o}M$.
 So another application of the dominated convergence theorem shows that
-${\mathrm{H}}^{F}(p)(w,w)=\int\limits_{M(\infty)}{\mathrm{H}}^{b_{v}}(p)(w,w)~{}\mathrm{d}\eta(v).$
+$\mathrm{Hess}\penalty 10000\ ^{F}(p)(w,w)=\int\limits_{M(\infty)}\mathrm{Hess}\penalty 10000\ b_{v}(p)(w,w)\penalty 10000\ \mathrm{d}\eta(v).$
 The convexity of $F$ follows from the convexity of the Busemann functions.
 We now claim that $F$ is, in fact, strictly convex.
-Suppose there exists $p\in M$ and $0\neq w\in T_{p}M$ such that ${\mathrm{H}}^{F}(p)(w,w)=0$ then
-$\int\limits_{M(\infty)}{\mathrm{H}}^{b_{v}}(p)(w,w)~{}\mathrm{d}\eta(v)=0$
-Since $p\mapsto{\mathrm{H}}^{b_{v}}(p)(w,w)$ is a nonnegative continuous function in $v$, we have ${\mathrm{H}}^{b_{v}}(p)(w,w)=0$ for all $v\in S_{o}M$. Using (2.2),
-(3.1) ${\langle D_{\nabla b_{v}(p)}^{{}^{\prime}}(0)w^{\perp}},{D_{\nabla b_{v}(p)}(0)w^{\perp}\rangle}=0,\quad\text{for all}~{}v\in S_{o}M.$
-It is known from [13, Section 5] that for $v\in SM$ and all parallel vector fields $x$ normal to $\gamma_{v}$, $\frac{d}{dt}(\left\lVert D_{v}(t)x\right\rVert^{2})\leq 0$ i.e., the function $f_{v,x}(t)=\frac{1}{2}\left\lVert D_{v}(t)x\right\rVert^{2}$ is $C^{2}$ and monotonically decreasing.
-From (3.1) we see that for each $v\in S_{o}M$, $f_{\nabla b_{v}(p),w^{\perp}}$ has a critical point at 00.
+Suppose there exists $p\in M$ and $0\neq w\in T_{p}M$ such that $\mathrm{Hess}\penalty 10000\ ^{F}(p)(w,w)=0$ then
+$\int\limits_{M(\infty)}\mathrm{Hess}\penalty 10000\ b_{v}(p)(w,w)\penalty 10000\ \mathrm{d}\eta(v)=0$
+Since $v\mapsto\mathrm{Hess}\penalty 10000\ b_{v}(p)(w,w)$ is a nonnegative continuous function in $v$, we have $\mathrm{Hess}\penalty 10000\ b_{v}(p)(w,w)=0$ for all $v\in S_{o}M$. Using (3),
+(6) ${\langle D_{\nabla b_{v}(p)}^{{}^{\prime}}(0)w^{\perp}},{D_{\nabla b_{v}(p)}(0)w^{\perp}\rangle}=0,\quad\text{for all}\penalty 10000\ v\in S_{o}M.$
+It is known from [Esc77, §5] that for $v\in SM$ and all parallel vector fields $x$ normal to $\gamma_{v}$, $\frac{d}{dt}(\left\lVert D_{v}(t)x\right\rVert^{2})\leq 0$ i.e., the function $f_{v,x}(t)=\frac{1}{2}\left\lVert D_{v}(t)x\right\rVert^{2}$ is $C^{2}$ and monotonically decreasing.
+From (6) we see that for each $v\in S_{o}M$, $f_{\nabla b_{v}(p),w^{\perp}}$ has a critical point at 0.
 We claim that $f_{\nabla b_{v}(p),w^{\perp}}^{{}^{\prime\prime}}(0)=0$.
-If not, then 00 is either a local minimum or local maximum.
-Suppose 00 is a local minimum then there exists a $\delta$-neighbourhood of 00 such that $f_{\nabla b_{v}(p),w^{\perp}}(0)\leq f_{\nabla b_{v}(p),w^{\perp}}(t)$ for all $t\in(-\delta,\delta)$.
+If not, then 0 is either a local minimum or local maximum.
+Suppose 0 is a local minimum then there exists a $\delta$-neighbourhood of 0 such that $f_{\nabla b_{v}(p),w^{\perp}}(0)\leq f_{\nabla b_{v}(p),w^{\perp}}(t)$ for all $t\in(-\delta,\delta)$.
 Since $f$ is monotonically decreasing we have $f_{\nabla b_{v}(p),w^{\perp}}(t)=f_{\nabla b_{v}(p),w^{\perp}}(0)$ for all $t\in(0,\delta)$.
@@
 Observe that $f_{\nabla b_{v}(p),w^{\perp}}$ is $C^{2}$ and so $f_{\nabla b_{v}(p),w^{\perp}}^{{}^{\prime\prime}}(0)=\lim\limits_{t\to 0_{+}}f_{\nabla b_{v}(p),w^{\perp}}^{{}^{\prime\prime}}(t)=0$, a contradiction.
-A similar argument could be carried out to show that 00 cannot be a local maximum of $f_{\nabla b_{v}(p),w^{\perp}}$.
-Hence 00 is a degenerate critical point of $f_{\nabla b_{v}(p),w^{\perp}}$ i.e, $f_{\nabla b_{v}(p),w^{\perp}}^{{}^{\prime\prime}}(0)=0$.
+A similar argument could be carried out to show that 0 cannot be a local maximum of $f_{\nabla b_{v}(p),w^{\perp}}$.
+Hence 0 is a degenerate critical point of $f_{\nabla b_{v}(p),w^{\perp}}$ i.e, $f_{\nabla b_{v}(p),w^{\perp}}^{{}^{\prime\prime}}(0)=0$.
 Then the Jacobi equation for $D_{\nabla b_{v}(p)}$,
@@
 gives
-00 0 $=f_{\nabla b_{v}(p),w^{\perp}}^{{}^{\prime\prime}}(0)$
+$0$ $=f_{\nabla b_{v}(p),w^{\perp}}^{{}^{\prime\prime}}(0)$
 $={\langle D_{\nabla b_{v}(p)}^{{}^{\prime\prime}}(0)w^{\perp}},{w^{\perp}\rangle}+\left\lVert D_{\nabla b_{v}(p)}^{{}^{\prime}}(0)w^{\perp}\right\rVert^{2}$
@@
 $\implies{\langle R_{\nabla b_{v}(p)}(0)w^{\perp}},{w^{\perp}\rangle}$ $=\left\lVert D_{\nabla b_{v}(p)}^{{}^{\prime}}(0)w^{\perp}\right\rVert^{2}$
-(3.2) $\implies K(\nabla b_{v}(p),w^{\perp})$ $=\left\lVert D_{\nabla b_{v}(p)}^{{}^{\prime}}(0)w^{\perp}\right\rVert^{2}\geq 0,\quad\text{for all $v\in S_{o}M$}$
+(7) $\implies K(\nabla b_{v}(p),w^{\perp})$ $=\left\lVert D_{\nabla b_{v}(p)}^{{}^{\prime}}(0)w^{\perp}\right\rVert^{2}\geq 0,\quad\text{for all $v\in S_{o}M$}$
 where $K$ denotes the sectional curvature corresponding to the plane spanned by $\nabla b_{v}(p)$ and $w^{\perp}$.
@@
 Thus the sequence $(w^{\perp}(m))$ converges to $w$ as $m\to\infty$ and from (4) we have,
-$K(\nabla b_{v_{k_{m}}}(p),w^{\perp}(m))\to K(x,w)\geq 0,\quad\text{as $m\to\infty$}$
+[…]
 Now for $\{x_{1},\cdots,x_{n-1}\}$ orthonormal vectors in $T_{p}M$ such that $x_{i}\perp w$, $Ricc(w,w)=\sum\limits_{i=1}^{n}K(x_{i},w)\geq 0$.
 This is a contradiction to the hypothesis that $M$ has negative Ricci curvature.
-Hence ${\mathrm{H}}^{F}(p)(w,w)>0$ for every $p\in M$ and every $0\neq w\in T_{p}M$ proving the strict convexity of $F$.
+Hence $\mathrm{Hess}\penalty 10000\ ^{F}(p)(w,w)>0$ for every $p\in M$ and every $0\neq w\in T_{p}M$ proving the strict convexity of $F$.
 ∎
-While there are several geometrically motivated convex functions on a Riemannian manifold, typically they have low regularity.
-Even in situations where such convex functions have higher regularity they often do not have other useful properties such as having bounded gradient.
-On the one hand, distance functions and Busemann functions are examples of convex functions with bounded gradient having low regularity.
-On the other hand, on complete simply connected non-compact Riemannian manifolds without focal points, the squared distance function is a smooth strictly convex function on $M$ [11, Lemma 1] whose gradient is not bounded.
-It is in this context that we have the following approximation result for strictly convex functions with bounded gradient.
-Note that the smoothing approximation fails if we drop the assumption of strict convexity [26].
-Recall that a $C^{2}$-function is strictly convex if its Hessian is positive definite.
-It was proved in [17, Theorem 2(a)] that a Lipschitz $C^{2}$-function which is strictly convex admits a smooth strictly convex approximation which is also Lipschitz.
-From this and using the fact that Lipschitz functions have bounded gradient we have the following corollary to our Theorem 3.1.
-Corollary 3.2.
-Let $M$ be as in Theorem 3.1. Then for any $C>1$, $M$ admits a smooth strictly convex function $f$ such that $\left\lVert\nabla f\right\rVert\leq C$.
-The property of having bounded gradient is useful in applications as can be seen below in Corollary 3.3 and Theorem 4.7.
-Let $(M,g,J)$ be a Kähler manifold as in the introduction.
-The complex analogue of convex functions is given by the notion of plurisubharmonic functions.
-These are defined using the complex Hessian called the Levi form.
-In local coordinates, the Levi form of a function $f$ is given by
-$Lf=4\sum\limits_{i,j=1}^{n}\frac{\partial^{2}f}{\partial z_{i}\partial\overline{z}_{j}}\mathrm{d}z_{i}\otimes\mathrm{d}\overline{z}_{j}$
-where $(z_{1},\cdots,z_{n})$ is a local holomorphic coordinate system.
-Let $(x_{1},y_{1},\cdots,x_{n},y_{n})$ be the corresponding normal coordinates with $z_{k}=x_{k}+iy_{k}$, $k=1,\cdots,n$.
-A real-valued function on a Kähler manifold is said to be (strictly) plurisubharmonic, if the Levi form is positive (definite) semi-definite.
-Since $M$ is Kähler, the Levi form $L_{f}$ of $f$ is expressed in terms of Hessian of $f$ using the real coordinates as
-$L_{f}={\mathrm{H}}^{f}\left(\frac{\partial}{\partial x_{1}},\frac{\partial}{\partial x_{1}}\right)+{\mathrm{H}}^{f}\left(\frac{\partial}{\partial y_{1}},\frac{\partial}{\partial y_{1}}\right).$
-Therefore if $f$ is a strictly convex $C^{2}$-function, then $f$ is strictly plurisubharmonic.
-As an application to Corollary 3.2, using [24, Lemma 4.2 (i)], we show the following volume growth comparison for metric balls.
-Corollary 3.3.
-If $M$ as in Theorem 3.1 is also Kähler of (complex) dimension $m$, then $M$ admits a smooth strictly plurisubharmonic function with bounded gradient.
-Moreover, for each $p\in M$ there exists a constant $C_{2}(p)>0$ such that ,
-$V_{p}(r)\geq C_{2}(p)r^{m},\quad r\geq 1$
-where $V_{p}(r)$ is the volume of a metric ball of radius $r$ about $p$ in $M$.
-4. Spectral theoretic applications
-We begin with a sketch of the basic functional analytic setup concerning the spectrum of the Laplacian following [4].
-Let $M$ be a complete Riemannian manifold.
-Then the (positive) Laplacian $\Delta$ is an essentially self-adjoint operator with the space of compactly supported smooth functions on $M$ as the domain.
-We also denote its unique self-adjoint extension, having the domain $\mathcal{D}(\Delta):=\{f\in L^{2}(M):\Delta f\in L^{2}(M)\}$, by $\Delta$.
-Using the functional calculus for $\Delta$ we have the family of spectral projection operators $\prod_{E}:=\chi_{E}(\Delta)$ for each Borel subset $E$ of ${\mathbb{R}}$ referred to as the resolution of $\Delta$.
-The support of these projections and the spectrum of $\Delta$ are related in the following way.
-We say that $\lambda\in{\mathbb{C}}$ is in the spectrum $\sigma(\Delta)$ of $\Delta$ if and only if $\prod_{(\lambda-\epsilon,\lambda+\epsilon)}\neq 0$ for all $\epsilon>0$.
-If $\prod_{\{\lambda\}}\neq 0$, then $\lambda$ is an eigenvalue of $\Delta$ and the range of $\prod_{\{\lambda\}}$ is the corresponding eigenspace.
-The spectral theorem gives a unitary equivalence between $\Delta$ and a multiplication operator on a suitable Hilbert space of the form $\mathrm{L}^{2}(X,\mu)$, where $\mu$ is a $\sigma-$ finite Borel measure on $X$.
-By the Lebesgue decomposition, $\mu$ decomposes as $\mu=\mu_{pp}+\mu_{ac}+\mu_{sc}$ into pure point measure, absolutely continuous and singularly continuous measures with respect to the Lebesgue measure, and consequently we have that
-(4.1) $\mathrm{L}^{2}(X,\mu)=\mathrm{L}^{2}(X,\mu_{pp})\oplus\mathrm{L}^{2}(X,\mu_{ac})\oplus\mathrm{L}^{2}(X,\mu_{sc}).$
-This induces a decomposition $\mathrm{L}^{2}(M)=\mathcal{H}_{pp}\oplus\mathcal{H}_{ac}\oplus\mathcal{H}_{sc}$ from which we have the associated decomposition of the spectrum of $\Delta$ namely, $\sigma(\Delta)=\overline{\sigma_{pt}(\Delta)}\cup\sigma_{ac}(\Delta)\cup\sigma_{sc}(\Delta)$ where
-$\sigma_{ac}(\Delta)=\sigma(\Delta_{|_{\mathcal{H}_{ac}}}),\quad\sigma_{sc}(\Delta)=\sigma(\Delta_{|_{\mathcal{H}_{sc}}}),\quad\overline{\sigma_{pt}(\Delta)}=\sigma(\Delta_{|_{\mathcal{H}_{pp}}})$
-Here we take the closure for the set of all eigenvalues as it is not necessarily closed.
-We refer to $\sigma_{ac}(\Delta)$ and $\sigma_{sc}(\Delta)$ as the absolutely continuous and singularly continuous parts of $\sigma(\Delta)$ respectively and their union as the continuous part $\sigma_{cont}(\Delta)$ of $\sigma(\Delta)$.
-In order to analyse the nature of the spectrum of the Laplacian, we invoke [28, Theorem 2].
-This result demands boundedness of the gradient and of the Laplacian of a strictly convex function as well as the additional constraint of nonpositivity on the bilaplacian and in turn yields the absolute continuity of the spectrum.
-Consequently, it enables us to exclude the existence of the point spectrum and the singularly continuous spectrum.
-As pointed out in [28], the seemingly unnatural additional condition that $\Delta^{2}F\leq 0$ is, in fact, necessary.
+As indicated at the beginning of the section, we will be concerned with the spectrum on asymptotically harmonic manifolds.
+Recall that these are simply connected Riemannian manifolds without conjugate points such that horospheres are of constant mean curvature $h\geq 0$.
+In particular, we answer more refined questions of whether there is a nontrivial point spectrum or singularly continuous spectrum on these manifolds.
+The main tool here is an implementation of Kato’s theory of $H$-smooth operators on complete Riemannian manifolds due to [Xav88, Theorem 2].
+When it can be carried out, it enables one to exclude the existence of the point spectrum and the singularly continuous spectrum.
+As noted in [Xav88], the seemingly unnatural additional condition $\Delta^{2}F\leq 0$ that is demanded is, in fact, necessary.
 Indeed, if this condition on the bilaplacian fails, Hadamard manifolds with eigenvalues in the spectrum are constructed there.
-The proof of absolute continuity of the spectrum is then derived from an estimate implementing Kato’s theory of $H-$smooth operators in Hilbert spaces.
 We shall observe below that the condition on the bilaplacian is satisfied for $M$ as in Theorem 4.1 which are in addition asymptotically harmonic.
-We first briefly discuss some facts regarding the bottom $\lambda_{0}(M)$ of the spectrum for asymptotically harmonic manifolds.
-For a complete noncompact Riemannian manifold $M$ it is characterized by [6, §4.4]
-(4.2) $\lambda_{0}(M)=\inf\limits_{f\neq 0}\frac{\int\limits_{M}|\nabla f|^{2}}{\int\limits_{M}|f|^{2}}$
-where $f$ varies over the completion $\mathfrak{H}(M)$ of $C_{c}^{\infty}(M)$ with respect to the norm $\left\lVert f\right\rVert=\left\lVert f\right\rVert_{\rm{L}^{2}(M)}+\left\lVert\nabla f\right\rVert_{\rm{L}^{2}(M)}$.
-Let $|\Omega|$ denote the volume of a submanifold $\Omega$ in $M$.
-Then the Cheeger’s constant of $M$ is given by [6, §4.3]
-$h_{Cheeger}(M):=\inf\limits_{\Omega\subset M}\frac{|\partial\Omega|}{|\Omega|}$
-where the infimum varies over all open, connected submanifolds $\Omega$ of $M$ with compact closure and smooth boundary $\partial\Omega$.
-The importance of this global geometric invariant is illustrated by the following fundamental inequality relating it to the global analytic invariant $\lambda_{0}(M)$ [6, §4.3, Theorem 3],
-(4.3) $\frac{h_{Cheeger}^{2}(M)}{4}\leq\lambda_{0}(M).$
-Proposition 4.1.
-Let $M$ be an asymptotically harmonic manifold with horospheres of mean curvature $h$ i.e. $\Delta b_{v}=h$ for all $v\in SM$. Then we have $h_{Cheeger}=h$ and $\lambda_{0}(M)=h^{2}/4$.
-Proof.
-For an open connected $\Omega$ in $M$ with compact closure and smooth boundary $\partial\Omega$ we have, by Stokes’ theorem, that
-$h|\Omega|=\int\limits_{\Omega}\Delta b_{v}=\left|\int\limits_{\partial\Omega}{\langle\nabla b_{v}},{\nu\rangle}\right|\leq|\partial\Omega|,\quad(\text{since$\left\lVert\nabla b_{v}\right\rVert=1$})$
-where $\nu$ is the outer unit normal vector field of $\Omega$ along $\partial\Omega$. Thus,
-(4.4) $h\leq\inf\limits_{\Omega\subseteq M}\frac{|\partial\Omega|}{|\Omega|}=h_{Cheeger}$
-where the infimum varies over $\Omega$ as above in $M$.
-Let $o\in M$ be arbitrary.
-Since $M$ is simply connected without conjugate points the exponential at $o$ is a diffeomorphism.
-Let $d(o,\cdot):M\to{\mathbb{R}}$ denote the distance from $o$.
-By taking the sequence of functions $f_{n}(x)=e^{-\frac{1}{2}(h+\frac{1}{n})d(o,x)}$ in $\mathfrak{H}(M)$, $n\in{\mathbb{N}}$, we see from (4.2) that
-(4.5) $\lambda_{0}(M)\leq\frac{h^{2}}{4}$
-Combining (4.3), (4.4), (4.5) we have
-(4.6) $h=h_{Cheeger}\quad\text{and}\quad\lambda_{0}(M)=\frac{h^{2}}{4}.$
-We remark here that for the proposition above to hold it suffices that there is a $v\in SM$ such that the corresponding horospheres have constant mean curvature.
 Lemma 4.2.
@@
 Since $\Delta$ is the trace of Hessian, by choosing an orthonormal basis $\{E_{i}(p)\}_{i=1}^{n}$ of $T_{p}M$ we have
-$\sum\limits_{i=1}^{n}{\mathrm{H}}^{F}(p)(E_{i}(p),E_{i}(p))$ $=\sum\limits_{i=1}^{n}\int\limits_{M(\infty)}{\mathrm{H}}^{b_{v}}(p)(E_{i}(p),E_{i}(p))~{}\mathrm{d}\eta(v)$
-$=\int\limits_{M(\infty)}\sum\limits_{i=1}^{n}{\mathrm{H}}^{b_{v}}(p)(E_{i}(p),E_{i}(p))~{}\mathrm{d}\eta(v)$
-$\Delta F(p)$ $=\int\limits_{M(\infty)}\Delta b_{v}(p)~{}\mathrm{d}\eta(v)=h$
+$\sum\limits_{i=1}^{n}\mathrm{Hess}\penalty 10000\ ^{F}(p)(E_{i}(p),E_{i}(p))$ $=\sum\limits_{i=1}^{n}\int\limits_{M(\infty)}\mathrm{Hess}\penalty 10000\ b_{v}(p)(E_{i}(p),E_{i}(p))\penalty 10000\ \mathrm{d}\eta(v)$
+$=\int\limits_{M(\infty)}\sum\limits_{i=1}^{n}\mathrm{Hess}\penalty 10000\ b_{v}(p)(E_{i}(p),E_{i}(p))\penalty 10000\ \mathrm{d}\eta(v)$
+$\Delta F(p)$ $=\int\limits_{M(\infty)}\Delta b_{v}(p)\penalty 10000\ \mathrm{d}\eta(v)=h$
 and the lemma follows. ∎
+As an immediate consequence we have the following theorem applying [Xav88, Theorem 2].
 Theorem 4.3.
-If $M$ as in Theorem 3.1 is also asymptotically harmonic with $h$ as the constant mean curvature of the horospheres, then the spectrum of the Laplacian is purely absolutely continuous and is contained in $[\frac{h^{2}}{4},\infty)$.
+If $M$ as in Theorem 4.1 is also asymptotically harmonic with $h$ as the constant mean curvature of the horospheres, then the spectrum of the Laplacian is purely absolutely continuous.
+Moreover, the spectrum is contained in $\frac{h^{2}}{4},\infty)$.
 Proof.
-The absolute continuity of the spectrum follows by combining [28, Theorem 2] with Lemma 4.2. That the spectrum of Laplacian is contained in $[\frac{h^{2}}{4},\infty)$ follows from (4.6). ∎
-Next we briefly describe harmonic manifolds, only mentioning aspects which are needed in the sequel and to set notation, and we refer to [1, Chapter $6$], and to the comprehensive exposition [21] for the details.
+The absolute continuity of the spectrum follows by combining [Xav88, Theorem 2] with Lemma 4.2.
+That the spectrum of Laplacian is contained in $\frac{h^{2}}{4},\infty)$ follows from the discussion below.
+∎
+We first briefly discuss some facts regarding the bottom $\lambda_{0}(M)$ of the spectrum for general asymptotically harmonic manifolds.
+For a complete noncompact Riemannian manifold $M$ it is characterized by [Cha84, §4.4]
+(8) $\lambda_{0}(M)=\inf\limits_{f\neq 0}\frac{\int\limits_{M}|\nabla f|^{2}}{\int\limits_{M}|f|^{2}}$
+where $f$ varies over the completion $\mathfrak{H}(M)$ of $C_{c}^{\infty}(M)$ with respect to the norm $\left\lVert f\right\rVert=\left\lVert f\right\rVert_{\rm{L}^{2}(M)}+\left\lVert\nabla f\right\rVert_{\rm{L}^{2}(M)}$.
+Let $|\Omega|$ denote the volume of a submanifold $\Omega$ in $M$.
+Then the Cheeger’s constant of $M$ is given by
+$h_{Cheeger}(M):=\inf\limits_{\Omega\subset M}\frac{|\partial\Omega|}{|\Omega|}$
+where the infimum varies over all open, connected submanifolds $\Omega$ of $M$ with compact closure and smooth boundary $\partial\Omega$.
+The importance of this global geometric invariant is illustrated by the following fundamental inequality relating it to the global analytic invariant $\lambda_{0}(M)$ [Cha84, §4.3, Theorem 3],
+(9) $\frac{h_{Cheeger}^{2}(M)}{4}\leq\lambda_{0}(M).$
+Proposition 4.4.
+Let $M$ be an asymptotically harmonic manifold with horospheres of mean curvature $h$ i.e., $\Delta b_{v}=h$ for all $v\in SM$. Then we have $h_{Cheeger}=h$ and $\lambda_{0}(M)=h^{2}/4$.
+Proof.
+For an open connected $\Omega$ in $M$ with compact closure and smooth boundary $\partial\Omega$ we have, by Stokes’ theorem, that
+$h|\Omega|=\int\limits_{\Omega}\Delta b_{v}=\left|\int\limits_{\partial\Omega}{\langle\nabla b_{v}},{\nu\rangle}\right|\leq|\partial\Omega|,\quad(\text{since $\left\lVert\nabla b_{v}\right\rVert=1$})$
+where $\nu$ is the outer unit normal vector field of $\Omega$ along $\partial\Omega$. Thus,
+(10) $h\leq\inf\limits_{\Omega\subseteq M}\frac{|\partial\Omega|}{|\Omega|}=h_{Cheeger}$
+where the infimum varies over $\Omega$ as above in $M$.
+Let $o\in M$ be arbitrary.
+Since $M$ is simply connected without conjugate points the exponential at $o$ is a diffeomorphism.
+Let $d(o,\cdot):M\to{\mathbb{R}}$ denote the distance from $o$.
+By taking the sequence of functions $f_{n}(x)=e^{-\frac{1}{2}(h+\frac{1}{n})d(o,x)}$ in $\mathfrak{H}(M)$, $n\in{\mathbb{N}}$, we see from (8) that
+(11) $\lambda_{0}(M)\leq\frac{h^{2}}{4}$
+Combining (9), (10), (11) we have
+(12) $h=h_{Cheeger}\quad\text{and}\quad\lambda_{0}(M)=\frac{h^{2}}{4}$
+Remark 4.5.
+We remark here that for the proposition above to hold it suffices that, in the case of simply connected Riemannian manifold without conjugate points, there is a single vector $v\in SM$ such that the corresponding horospheres have constant mean curvature.
+We have the following immediate corollary to Theorem 4.3 following [Don81] when the sectional curvature is strictly negative.
+Corollary 4.6.
+Let $M$ be an asymptotically harmonic Hadamard manifold with strictly negative sectional curvature.
+Then the spectrum of the Laplacian is purely absolutely continuous and is equal to $\frac{h^{2}}{4},\infty)$.
+Observe that for such manifolds the essential spectrum and the absolutely continuous spectrum have been determined in [CS14, Theorem 4.4] and [BMP23, Theorem 1.3] respectively.
+Harmonic manifolds are important examples of asymptotically harmonic manifolds.
+We describe only the aspects of harmonic manifolds which are needed in the sequel and to set notation, and we refer to [Bes78, Chapter $6$], and to the comprehensive exposition [Kni16] for the details.
 A function $f$ on a Riemannian manifold $(M,g)$ is said to be radial around a point $o$ if $f$ is constant on geodesic spheres $S_{o}(r)$ of radius $r$ centered at $o$.
@@
     If $\kappa=0$, then $M$ is Euclidean.
-Thus simply connected noncompact nonflat harmonic manifolds necessarily have constant negative Ricci curvature.
-For every $v\in SM$, the Busemann function $b_{v}$ satisfies $\Delta b_{v}=h$ (see for instance [5, §2, p.
+Thus simply connected noncompact nonflat harmonic manifolds necessarily have constant negative Ricci curvature $\kappa<0$.
+For every $v\in SM$, the Busemann function $b_{v}$ satisfies $\Delta b_{v}=h$ (see for instance [Bra24, §2, p.
 33]) and so it follows that every harmonic manifold is also asymptotically harmonic.
@@
 Corollary 4.7.
-Let $M$ be a complete simply connected noncompact nonflat harmonic manifold $M$ without focal points.
-Then the spectrum of the Laplacian is purely absolutely continuous and is contained in $[h^{2}/4,\infty)$.
-We now proceed to show that the spectrum is actually equal to the interval $[h^{2}/4,\infty)$, by making use of the unitarity of the spherical Fourier transform for radial functions on harmonic manifolds of purely exponential volume growth studied in [3].
-Notice that in [3] unitarity is shown only in dimensions greater than $5$.
-However, the only simply connected noncompact nonflat harmonic manifolds of dimension less than equal to $5$ are Riemannian symmetric spaces of noncompact type.
-This case is handled in Theorem 4.8 below.
-First, we observe from [21, Theorems 32, 45] that, noncompact harmonic manifolds without focal points fall into the class of noncompact harmonic manifolds of purely exponential volume growth.
+Let $M$ be a simply connected noncompact (nonflat) harmonic manifold $M$ without focal points. Then the spectrum of the Laplacian is purely absolutely continuous and is contained in $h^{2}/4,\infty)$.
+Remark 4.8.
+We reiterate here that for this class of manifolds, the negativity of Ricci curvature is not an additional assumption.
+We now proceed to give a novel proof that the spectrum is actually equal to the interval $h^{2}/4,\infty)$.
+For this, we make use of the recently proved [BKP21] unitarity of the Fourier transform for radial functions on harmonic manifolds of purely exponential volume growth, certain estimates on the $c$-function, and the characterisation of harmonic manifolds that the radialisation operator commutes with the Laplacian to pass from radial $\mathrm{L}^{2}$-functions to all $\mathrm{L}^{2}$-functions.
+We note that unitarity there is shown only in dimensions greater than $5$.
+However, the only simply connected noncompact (nonflat) harmonic manifolds of dimension less than equal to $5$ are Riemannian symmetric spaces of noncompact type [BKP21, p.145] and this case is covered by the discussion on symmetric spaces at the end.
+First, we observe from [Kni16, Theorems 32, 45] that, noncompact harmonic manifolds without focal points fall into the class of noncompact harmonic manifolds of purely exponential volume growth.
 Let $o$ be an arbitrary but fixed point in $M$.
@@
 for $\lambda\in{\mathbb{C}}$ and $v\in S_{o}M$.
-Here $\phi_{\lambda}$ is the unique function on $[0,\infty)$ with $\phi_{\lambda}(0)=1$ and satisfying $\Delta(\phi_{\lambda}\circ d_{o})=(\lambda^{2}+\frac{h^{2}}{4})(\phi_{\lambda}\circ d_{o})$, where $d_{o}$ denotes the distance from $o$.
-Let ${\rm{L}}_{o}^{2}(M)$ denote the closed subspace of ${\rm{L}}^{2}(M)$ consisting of those functions which are radial around $o$ in ${\rm{L}}^{2}(M)$.
-It is known from [3, Theorem 4.7] that the ${\mathcal{F}}^{o}$ is an unitary operator from ${\rm{L}}_{o}^{2}(M)$ onto ${\rm{L}}^{2}([0,\infty),C_{0}|c(\lambda)|^{-2}{\rm{d}}\lambda)$.
+Here $\phi_{\lambda}$ is the unique function on $0,\infty)$ with $\phi_{\lambda}(0)=1$ and satisfying $\Delta(\phi_{\lambda}\circ d_{o})=(\lambda^{2}+\frac{h^{2}}{4})(\phi_{\lambda}\circ d_{o})$, where $d_{o}$ denotes the distance from $o$.
+Let ${\rm{L}}_{o}^{2}(M)$ denote the closed subspace of ${\rm{L}}^{2}(M)$ consisting of those functions which are radial about $o$ in ${\rm{L}}^{2}(M)$.
+It is known from [BKP21, Theorem 4.7] that ${\mathcal{F}}^{o}$ is a unitary operator from ${\rm{L}}_{o}^{2}(M)$ onto ${\rm{L}}^{2}([0,\infty),C_{0}|c(\lambda)|^{-2}{\rm{d}}\lambda)$.
 Here the $c$-function coincides with the Harish Chandra $c$-function on $Im(\lambda)\leq 0$, for rank one symmetric spaces of noncompact type.
 We recall that for a real-valued measurable function $f$ on a $\sigma$-finite measure space $(X,\mu)$, the multiplication operator $M_{f}$ on ${\rm{L}}^{2}(X,\mu)$ is self-adjoint with domain $\mathcal{D}(M_{f}):=\{g\in{\rm{L}}^{2}(X,\mu):M_{f}g\in{\rm{L}}^{2}(X,\mu)\}$.
-Let $m$ be the function on $[0,\infty)$ defined by $m(\lambda)=\lambda^{2}+\frac{h^{2}}{4}$.
+Let $m$ be the function on $0,\infty)$ defined by $m(\lambda)=\lambda^{2}+\frac{h^{2}}{4}$.
 Then we see that the multiplication operator $M_{m}$ on ${\rm{L}}^{2}([0,\infty),C_{0}|c(\lambda)|^{-2}{\rm{d}}\lambda)$ is self-adjoint on
 $\mathcal{D}(M_{m})=\{f\in{\rm{L}}^{2}([0,\infty),C_{0}|c(\lambda)|^{-2}{\rm{d}}\lambda):M_{m}f\in{\rm{L}}^{2}([0,\infty),C_{0}|c(\lambda)|^{-2}{\rm{d}}\lambda)\}.$
-Moreover, the spectrum of a multiplication operator $M_{m}$ is equal to the essential range of $m$ [4, Theorem 4.5]. Here
-$ess-range(m):=\{x\in{\mathbb{R}}:\mu(m^{-1}(B(x,\epsilon)))>0~{}\forall~{}\epsilon>0\}.$
-Now to determine the spectrum of $M_{m}$ we argue as follows. For any $x\in{\mathbb{R}}$ and $\epsilon>0$,
-$m^{-1}(B(x,\epsilon))=\{\lambda\in[0,\infty):x-\epsilon<\lambda^{2}+\frac{h^{2}}{4}<x+\epsilon\}$
-Note that when $x<\frac{h^{2}}{4}$ and $\epsilon=-(x-\frac{h^{2}}{4})$, $m^{-1}(B(x,\epsilon))$ is empty and so $x\notin ess-range(m)$.
+Moreover, the spectrum of a multiplication operator $M_{m}$ is equal to the essential range of $m$ [Bor20, Theorem 4.5]. Here
+$ess-range(m):=\{x\in{\mathbb{R}}:|m^{-1}(x-\epsilon,x+\epsilon))|>0,\penalty 10000\ \forall\penalty 10000\ \epsilon>0\}.$
+where $|\cdot|$ denotes the measure. Now to determine the spectrum of $M_{m}$ we argue as follows. For any $x\in{\mathbb{R}}$ and $\epsilon>0$,
+$m^{-1}(x-\epsilon,x+\epsilon)=\{\lambda\in[0,\infty):x-\epsilon<\lambda^{2}+\frac{h^{2}}{4}<x+\epsilon\}$
+Note that when $x<\frac{h^{2}}{4}$ and $\epsilon=-(x-\frac{h^{2}}{4})$, $m^{-1}(x-\epsilon,x+\epsilon)$ is empty and so $x\notin ess-range(m)$.
 Equivalently, $ess-range(m)\subseteq[\frac{h^{2}}{4},\infty)$.
 Now for any $x\in[\frac{h^{2}}{4},\infty)$ and $\epsilon>0$ we have,
-$m^{-1}(B(x,\epsilon))=\begin{cases}\left[0,\sqrt{x-\frac{h^{2}}{4}+\epsilon}\right)&\text{if}~{}x-\frac{h^{2}}{4}-\epsilon<0\\ \\ \left[\sqrt{x-\frac{h^{2}}{4}-\epsilon},\sqrt{x-\frac{h^{2}}{4}+\epsilon}\right)&\text{if}~{}x-\frac{h^{2}}{4}-\epsilon\geq 0.\\ \end{cases}$
-Using the estimates on $|c(\lambda)|^{-1}$ given in [3, §4.1] we see that, for $C>0$ and $K=\sqrt{x-\frac{h^{2}}{4}+\epsilon}>0$,
-$\int\limits_{m^{-1}(B(x,\epsilon))}C_{0}|c(\lambda)|^{-2}{\rm{d}}\lambda\geq\frac{C_{0}}{C}\int\limits_{m^{-1}(B(x,\epsilon))}|\lambda|^{2}{\rm{d}}\lambda>0.$
-It follows that $ess-range(m)=[\frac{h^{2}}{4},\infty)$ and so the spectrum of $M_{m}$ is equal to $[\frac{h^{2}}{4},\infty)$.
-For functions that are not necessarily radial, a notion of Fourier transform, relative to a point, was introduced in [3, §5] where it was shown to be an isometry but the unitarity is not known.
-This Fourier transform coincides with the spherical Fourier transform ${\mathcal{F}}^{o}$ for radial functions [3, Proposition 5.3].
-Then by using [5, Lemma 6.5], we see that
+[…]
+We recall the estimates for $|c(\lambda)|^{-1}$ given in [BKP21, §4.1].
+If $x-\frac{h^{2}}{4}-\epsilon<0$, then $\lambda\in\left[0,\sqrt{x-\frac{h^{2}}{4}+\epsilon}\right)$. If $\sqrt{x-\frac{h^{2}}{4}+\epsilon}\leq K$ then we have
+$\frac{1}{C}\lambda\leq|c(\lambda)|^{-1}$
+and so
+$\int\limits_{m^{-1}(x-\epsilon,x+\epsilon)}C_{0}|c(\lambda)|^{-2}{\rm{d}}\lambda\geq\frac{C_{0}}{C}\int\limits_{0}^{\sqrt{x-\frac{h^{2}}{4}+\epsilon}}\lambda^{2}{\rm{d}}\lambda>0,\quad\text{since}\penalty 10000\ x-\frac{h^{2}}{4}+\epsilon>0$
+Now, if $K\in\left(0,\sqrt{x-\frac{h^{2}}{4}+\epsilon}\right)$ then
+$\begin{cases}\frac{1}{C}\lambda\leq|c(\lambda)|^{-1},&\lambda\in[0,K)\\ \frac{1}{C}\lambda^{\frac{n-1}{2}}\leq|c(\lambda)|^{-1},&\lambda\in\left[K,\sqrt{x-\frac{h^{2}}{4}+\epsilon}\right)\end{cases}$
+holds thereby
+$\int\limits_{m^{-1}(x-\epsilon,x+\epsilon)}C_{0}|c(\lambda)|^{-2}{\rm{d}}\lambda\geq\begin{cases}\frac{C_{0}}{C}\int\limits_{0}^{K}\lambda^{2}{\rm{d}}\lambda,&\lambda\in[0,K)\\ \\ \frac{C_{0}}{C}\int\limits_{K}^{\sqrt{x-\frac{h^{2}}{4}+\epsilon}}\lambda^{n-1}{\rm{d}}\lambda,&\lambda\in\left[K,\sqrt{x-\frac{h^{2}}{4}+\epsilon}\right)\end{cases}$
+and hence the integral is always strictly positive.
+A similar argument shows that the integral remains strictly positive for the other case in (13).
+It follows that $ess-range(m)=[\frac{h^{2}}{4},\infty)$ and so the spectrum of $M_{m}$ is equal to $\frac{h^{2}}{4},\infty)$.
+For functions that are not necessarily radial, a notion of Fourier transform, relative to a point, was introduced in [BKP21, §5] where it was shown to be an isometry but the unitarity is not known.
+This Fourier transform coincides with the spherical Fourier transform ${\mathcal{F}}^{o}$ for radial functions [BKP21, Proposition 5.3].
+Then we see as in [Bra24, Lemma 6.5], that
 ${\mathcal{F}}^{o}\Delta f=M_{m}{\mathcal{F}}^{o}f,\quad f\in{\rm{L}}_{o}^{2}(M).$
@@
 Lemma 4.9.
-The spectrum of the Laplacian $\Delta$ on $\mathcal{D}(\Delta)\cap{\rm{L}}_{o}^{2}(M)$ is equal to $[h^{2}/4,\infty)$.
+The spectrum of the Laplacian $\Delta$ on $\mathcal{D}(\Delta)\cap{\rm{L}}_{o}^{2}(M)$ is equal to $h^{2}/4,\infty)$.
 We now claim that,
 Lemma 4.10.
-$[h^{2}/4,\infty)\subseteq\sigma(\Delta)$ where $\sigma(\Delta)$ is the spectrum of $\Delta$ with domain $\mathcal{D}(\Delta)$.
+$h^{2}/4,\infty)\subseteq\sigma(\Delta)$ where $\sigma(\Delta)$ is the spectrum of $\Delta$ with domain $\mathcal{D}(\Delta)$.
 Proof.
@@
 $(\Delta-\mu)g=f=R_{o}f=R_{o}(\Delta-\mu)g=(\Delta-\mu)R_{o}g$
-and by uniqueness $R_{o}g=g$ i.e., $\Delta-\mu$ is surjective from $\mathcal{D}(\Delta)\cap{\rm{L}}_{o}^{2}(M)$ onto ${\rm{L}}_{o}^{2}(M)$ and hence invertible.
+and since $\Delta-\mu$ is invertible we have $R_{o}g=g$ i.e., $\Delta-\mu$ is surjective from $\mathcal{D}(\Delta)\cap{\rm{L}}_{o}^{2}(M)$ onto ${\rm{L}}_{o}^{2}(M)$.
 Since $(\Delta-\mu)^{-1}$ is bounded on ${\rm{L}}^{2}(M)$ and ${\rm{L}}_{o}^{2}(M)$ is a closed subspace, $(\Delta-\mu)^{-1}$ is also bounded on $\mathcal{D}(\Delta)\cap{\rm{L}}_{o}^{2}(M)$.
@@
 ∎
-The equality of the spectrum now follows from Corollary 4.4 and Lemma 4.6. Therefore, we have the following theorem.
-Theorem 4.7.
-Let $M$ be a simply connected noncompact nonflat harmonic manifold without focal points. Then the spectrum of Laplacian is purely absolutely continuous and is equal to $[h^{2}/4,\infty)$.
-[…]
+The equality of the spectrum is obtained from Corollary 4.7 and Lemma 4.10. In summary, we have the following proposition.
+Proposition 4.11.
+Let $M$ be a simply connected noncompact (nonflat) harmonic manifold without focal points. Then the spectrum of Laplacian is purely absolutely continuous and is equal to $h^{2}/4,\infty)$.
+As remarked before, rank-one symmetric spaces of noncompact type, Damek-Ricci spaces are among the classes of manifolds which satisfy the hypothesis of Proposition 4.11.
+In fact, it is known that an irreducible symmetric space is harmonic if and only if it is rank one [Kni16].
+We shall see that, on the other hand, Riemannian symmetric spaces of noncompact type and higher rank are not even asymptotically harmonic and therefore Theorem 4.3 does not apply.
+Nonetheless, we will indicate below that our method can still be employed to show an analogue of Proposition 4.11 in this setting.
+Let $M$ be an $n$-dimensional symmetric space of noncompact type i.e., $M$ is a simply connected Riemannian manifold with parallel curvature tensor $R$ [Ebe96, Chapter 2].
+Note that $M$ is necessarily of the form $G/K$ where $G=NAK$ is the Iwasawa decomposition of a noncompact connected semisimple Lie group $G$ with finite center.
+The rank of the symmetric space is the dimension $d$ of $\mathfrak{a}$, the Lie algebra of $A$.
+The Killing form on the Lie algebra of $G$ induces an inner product on $\mathfrak{a}$ denoted by $\langle\cdot$$\cdot\rangle$}.
+Let us fix a base point $o=eK\in G/K$.
+If the rank of $M$ is atleast two, then its sectional curvature vanishes on the $G$-conjugates of $A.o$ in $G/K$.
+These are the maximal totally geodesic flat submanifolds in $G/K$.
+For $v\in S_{o}M$, consider the geodesic $\gamma_{v}$ satisfying $\gamma_{v}^{{}^{\prime}}(0)=v$ and let $R_{v}$ be the curvature tensor restricted to $\gamma_{v}$.
+Choose an orthonormal basis $\{e_{1},\cdots,e_{n-1}\}$ of eigenvectors of $R_{v}$ for the subspace in $T_{o}M$ orthogonal to $v$ i.e.,
 $R_{v}e_{i}=-R(v,e_{i})v=\lambda_{i}(v)e_{i}.$
@@
 We denote the tangent vector that is dual to $2\rho=\sum\limits_{\alpha\in\sum^{+}}m_{\alpha}\alpha$ also by $2\rho$.
-The arguments in [27, Lemma 1] show that for each $v\in S_{o}M$, $\Delta b_{v}$ is the constant function given by $\sum\limits_{i}\sqrt{\lambda_{i}(v)}={\langle v},{2\rho\rangle}$.
+The arguments in [Wan15, Lemma 1] show that for each $v\in S_{o}M$, $\Delta b_{v}$ is the constant function given by $\sum\limits_{i}\sqrt{\lambda_{i}(v)}={\langle v},{2\rho\rangle}$.
 This constant is the mean curvature of horospheres with respect to $b_{v}$.
@@
 In particular, for $v=\frac{\rho}{\left\lVert\rho\right\rVert}$ we have $\left\lVert\rho\right\rVert\leq h_{Cheeger}$ and hence by Cheeger’s inequality we obtain
-(4.7) $\left\lVert\rho\right\rVert^{2}\leq\lambda_{0}(M).$
-Observe that $M$ has negative Ricci curvature [12, §2.14, p. 96] and has no focal points. With notations as in Theorem 3.1 we have,
+(14) $\left\lVert\rho\right\rVert^{2}\leq\lambda_{0}(M).$
+Observe from [Wan15], that along a geodesic $\gamma_{v}$ the Jacobi field $Y_{i}$ satisfying the initial condition $Y_{i}(0)=0$ and $Y_{i}^{{}^{\prime}}(0)=e_{i}$ is given by the explicit formula
+$Y_{i}(t)=\frac{1}{\sqrt{\lambda_{i}(v)}}\sinh\left(\sqrt{\lambda_{i}(v)}t\right)E_{i}(t)$
+where $E_{i}$ is a parallel orthonormal frame field along $\gamma_{v}$ with $E_{i}(0)=e_{i}$. Then we see that
+$Y(t)=\sum\limits_{i=1}^{n}{\langle w},{e_{i}\rangle}Y_{i}(t)$
+solves the Jacobi equation (1) with initial conditions $Y(0)=0$ and $Y^{{}^{\prime}}(0)=w$ for arbitrary $w\in SM$.
+That $\left\lVert Y(t)\right\rVert$ is strictly increasing can then be deduced from the fact that $\sinh$ is an increasing function and so, it follows from [O’S74, Proposition 4] that $M$ has no focal points.
+It is known that $M$ is also Einstein with negative Ricci curvature [Ebe96, §2.14, p.
+96].
+Therefore Theorem 4.1 applies and we have,
   1. (i)
@@
     For each $v\in S_{o}M$, $\Delta b_{v}$ is the constant function ${\langle v},{2\rho\rangle}$ and so we have $\Delta^{2}b_{v}=0$. It follows that, $\Delta^{2}F=0$.
-The absolute continuity of the spectrum $\sigma(\Delta)$ is then established by invoking [28, Theorem 2]. Further, using (4.7) we have,
-Theorem 4.8.
-Let $M$ be a symmetric space of noncompact type. Then the spectrum of the Laplacian on $M$ is absolutely continuous and is equal to $[\left\lVert\rho^{2}\right\rVert,\infty)$.
-Here we use the spherical Fourier transform for the $K$-bi-invariant functions on $G$ to prove the equality of the spectrum.
-The radial part of the Laplacian $\Delta$ for the action of $K$ on $G/K$ is given by [19, §4.5, p.
-427],
-(4.8) $\Delta_{rad}\phi_{\lambda}=({\langle\lambda},{\lambda\rangle}+{\langle\rho},{\rho\rangle})\phi_{\lambda},\quad\lambda\in\mathfrak{a}_{{\mathbb{C}}}^{\ast}$
-where $\mathfrak{a}_{{\mathbb{C}}}^{\ast}$ the complexification of the dual $\mathfrak{a}^{\ast}$ of $\mathfrak{a}$ and $\phi_{\lambda}(gK)=\displaystyle{\int\limits_{K}e^{(i\lambda+\rho)A(kg)}~{}{\rm{d}}k}$ is the unique spherical function on $G$ with the normalization $\phi_{\lambda}(eK)=1$ and $A$ is referred to as the Iwasawa projection $A:G\to{\mathfrak{a}}$.
-Two spherical functions are identical up to the action by the Weyl group $W$ on $\mathfrak{a}_{{\mathbb{C}}}^{\ast}$.
-Let us denote the space of smooth compactly supported $K-$invariant functions on $G/K$ by $C_{c}^{\infty}(K\backslash G/K)$ and its closure in ${\rm{L}}^{2}(G/K)$ by ${\rm{L}}_{rad}^{2}(G/K)$.
-Using the inversion formula for the spherical transform [19, Chapter 4, Theorem 7.5] the map $f\mapsto\widehat{f}$
-$\widehat{f}(\lambda)=\int\limits_{G/K}f(gK)\phi_{\lambda}(gK),\quad f\in C_{c}^{\infty}(K\backslash G/K)$
-[…]
-(4.9) $\int\limits_{m^{-1}(B(x,\epsilon))}|\textbf{c}(\lambda)|^{-2}{\rm{d}}\lambda>0.$
-for $x\geq\left\lVert\rho\right\rVert^{2}$ and $\epsilon>0$.
-Note that $|\textbf{c}(\lambda)|^{2}$ defined on ${\mathfrak{a^{\ast}}}$ is invariant under the action of $W$ i.e., $|\textbf{c}(w\lambda)|^{2}=|\textbf{c}(\lambda)|^{2}$ for all $w\in W$ [19, §IV.7, Equation 16] and hence descends to a function on ${\mathfrak{a}^{\ast}}/W$.
-Since $\lambda\mapsto\left\lVert\lambda\right\rVert^{2}$ is continuous and
-$m^{-1}(B(x,\epsilon))=\left\{\lambda\in{\mathfrak{a}^{\ast}}/W:x-\epsilon-\left\lVert\rho\right\rVert^{2}<\left\lVert\lambda\right\rVert^{2}<x+\epsilon-\left\lVert\rho^{2}\right\rVert\right\}$
-we see that $m^{-1}(B(x,\epsilon))$ contains an open ball in $\mathfrak{a}^{\ast}/W$.
-Using [15, Chapter 3, p.
-101, Theorem 6.3.4], we have that the map $\lambda\mapsto|\textbf{c}(\lambda)|^{-2}$ is analytic on $\mathfrak{a}^{\ast}$ and thereby analytic on $\mathfrak{a}^{\ast}/W$.
-Thus, (4.9) holds for $x\geq\left\lVert\rho\right\rVert^{2}$ and $\epsilon>0$.
-This, in particular, shows that the spectrum of $\Delta_{rad}$ is equal to $[\left\lVert\rho\right\rVert^{2},\infty)$.
-We argue as in Lemma 4.6 to show the equality in Theorem 4.8.
+As before, the absolute continuity of the spectrum $\sigma(\Delta)$ is established by using [Xav88, Theorem 2].
+That the spectrum is equal to $\left\lVert\rho^{2}\right\rVert,\infty)$ can be shown analogous to the case of harmonic manifolds by using the spherical Fourier transform and estimates on the c-function found in [GV88, Chapter 3, p.
+101, Theorem 6.3.4].
 5. Funding